### v4

# Do Language Models Associate Sound with Meaning? A Multimodal Study of Sound Symbolism

Abstract
Sound symbolism is a linguistic concept that refers to non-arbitrary associations between phonetic forms and their meanings. We suggest that this can be a compelling probe into how Multimodal Large Language Models (MLLMs) interpret auditory information in human languages. We investigate MLLMs' performance on phonetic iconicity across textual (orthographic and IPA) and auditory forms of inputs with up to 25 semantic dimensions (e.g., sharp vs. round), observing models' layer-wise information processing by measuring phoneme-level attention fraction scores. To this end, we present LEX-ICON, an extensive mimetic word dataset consisting of 8,052 words from four natural languages (English, French, Japanese, and Korean) and 2,930 systematically constructed pseudo-words, annotated with semantic features applied across both text and audio modalities. Our key findings demonstrate (1) MLLMs' phonetic intuitions that align with existing linguistic research across multiple semantic dimensions and (2) phonosemantic attention patterns that highlight models' focus on iconic phonemes. These results bridge domains of artificial intelligence and cognitive linguistics, providing the first large-scale, quantitative analyses of phonetic iconicity in terms of MLLMs' interpretability.
Codehttps://github.com/jjhsnail0822/sound-symbolism
1 Introduction
Figure 1: Phonetic iconicity investigation for MLLMs using natural and constructed mimetic words from text and audio modalities in LEX-ICON. We conduct quantitative evaluations for up to 25 semantic dimensions and examine layer-wise attention fraction scores to identify how phonemes and meanings are related within the models.
Sound symbolism, which suggests that phonetic sounds and their meanings have a significant correlation, presents a cognitively grounded exception to the principle of linguistic arbitrariness (hinton2006sound; dingemanse2015arbitrariness). For instance, when people are presented with two images of sharp and round shapes called ``kiki'' and ``bouba'', the overwhelming majority of participants consistently match ``kiki'' with pointed shapes and ``bouba'' with round shapes, regardless of their cultural background (kohler1967gestalt; ramachandran2001synaesthesia). This iconicity effect illustrates the universal sensory associations of humans that intuitively link certain phonetic sounds with certain semantic features, which facilitate early-childhood or second-language acquisition (imai2014sound), and commercially nuanced brand naming (yorkston2004sound).
Recent advancements in Multimodal Large Language Models (MLLMs) that incorporate audio-modality input into an integrated representation space (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport) shed light on new possibilities to systematically analyze human-like phonetic iconicity of language models. With MLLMs as a novel test bed, we formulate the following two key research questions:
  1. RQ 1.
    How do MLLMs associate sound-symbolic words with semantic features similar to human phonetic intuition?
  2. RQ 2.
    How do MLLMs' internal attention patterns align with phonosemantic relationships?
Mimetic words provide a compelling probe to address these subjects. Represented by onomatopoeias and ideophones [^1] In this paper, “onomatopoeia” and “ideophone” refer to words that depict sounds and non-auditory sensory imagery, respectively., mimetic words refer to words in which the non-arbitrary association of sound form and meaning is prominent (akita2019ideophones), such as ``boom'' or ``whizz.'' We construct LEX-ICON, a dataset of natural mimetic words and constructed pseudo-words designed to maximize sound-symbolic effect. With LEX-ICON, we apply the semantic dimension paradigm, where word meanings are projected onto binary scales (e.g., fast vs. slow), to examine MLLMs' abilities to infer the meaning of words from their form (original text, IPA, audio) rather than their content. We further analyze models' internal representations to identify whether models actually focus on iconic phonemes, as illustrated in Figure 1.
Our experimental results show that MLLMs demonstrate phonetic intuitions across multiple semantic dimensions, for not only natural words but also constructed pseudo-words that exclude models' memorization (e.g., sharp vs. round). In particular, MLLMs exhibit modality-specific preferences across semantic dimensions, using audio for acoustically grounded features (e.g., big vs. small) and text for articulatory or visually driven features (e.g., beautiful vs. ugly). However, we reveal that there still remain discrepancies between the semantic dimensions where humans exhibit high scores and those where models perform well. Through phoneme-level attention analysis, we suggest that MLLMs attend to sound-symbolic phonemes corresponding to their meanings. In particular, we demonstrate that attention fraction to these sound-symbolic phonemes is more prominent in models' late-layers when processing constructed pseudo-words, suggesting that the relationship between phonemes and meanings can be systematically represented in the deep layers. Based on these findings, we summarize our key contributions as follows.
  1. 1.
    We conduct the first large-scale investigation of phonetic iconicity in MLLMs, grounded in LEX-ICON, a novel multilingual mimetic word dataset spanning multiple language families.
  2. 2.
    We provide a quantitative evaluation methodology of sound symbolism through the semantic dimension approach that contributes to linguistics by capturing phonetic iconicity akin to humans.
  3. 3.
    We present a comprehensive analysis that elucidates models' internal mechanisms towards sound symbolism by measuring phoneme-level attention scores, thereby contributing to the field of model interpretability.
Our approach integrates two distinct but interconnected domains, addressing both artificial intelligence and linguistics. For the former, we observe an integration of form and meaning within MLLMs using sound symbolism probes, revealing new insights in terms of model interpretability. From a linguistic perspective, we propose a substantiation for nonhuman intelligence's phonetic intuitions on mimetic word data, providing a quantitative basis for linguistic experiments that have been conducted mostly with humans.
2 Related Work
2.1 Sound Symbolism in Linguistics
Early modern linguists question the arbitrary relationship between signifier and signified, suggesting the phenomenon of intuitive phonetic symbolism in humans (usnadze1924experimenteller; sapir1929study; kohler1967gestalt). sapir1929study's experiment has exhibited that people who are presented with unfamiliar object names ``mil'' and ``mal'' and are asked to estimate the size of the objects respond that the latter is larger than the former (parise2012audiovisual). Recent studies continue to observe these phenomena (hinton2006sound; lockwood2015iconicity; cwiek2022bouba), extending them by scaling to a large amount of data (thompson2021articulatory; winter2024iconicity) or measuring the exact semantic dimensions associated with each phoneme (monaghan2019sound; sidhu2022higher; sidhu2025sound). For the mimetic words, experimental research has also demonstrated that people can infer a certain degree of meaning from the words in languages besides their mother tongue (shinohara2010cross; dingemanse2016sound).
2.2 Phonetic Iconicity for LLMs
Iconicity tasks for language models have been used as a means of assessing whether the models have human-like phonetic intuition (cai-etal-2024-large; duan2024hlbbenchmarkingllmshumanlikeness). Early studies focus on analyzing phonesthemic information contained in word embedding spaces (abramova2013automatic; abramova-fernandez-2016-questioning). Recent studies demonstrate that the effect of non-arbitrary integration of linguistic form and meaning is substantiated in text-only LLMs (miyakawa-etal-2024-llms; marklova2025iconicitylargelanguagemodels), vision and image generation models (loakman-etal-2024-ears; alper2024kikiboubasoundsymbolism; shinto2024analyzing; iida2024investigating), and audio-visual models (tseng2024measuring), yet they are not extensive enough in scope to include exhaustive and multilingual mimetic word data or diverse semantic feature analysis.
2.3 Multimodal Interpretability
Recently, the field of model interpretability has witnessed significant growth (elhage2021mathematical; zou2025representationengineeringtopdownapproach; wang2023interpretability). Motivated by the rapid progress of MLLMs (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport), there has been increasing interest in multimodal interpretability (lin2025surveymechanisticinterpretabilitymultimodal). For instance, neo2025towards ablates a subset of visual tokens to observe the resulting differences in model output, while nikankin2025taskdifferentcircuitsdisentangling demonstrates that models employ distinct processing circuits depending on the input modality. Despite these advances, prior work has predominantly focused on the visual modality, with less exploration of audio-based interpretability. yang2025audiolenscloserlookauditory investigates how models handle auditory input, but their analysis is limited to simple settings where the audio simply serves as a direct vocalization of textual data. In this work, we examine audio interpretability through the lens of sound symbolism, highlighting how phoneme-level features encode meaning in ways that are intrinsically tied to the auditory modality.
3 LEX-ICON
Figure 2: A comprehensive figure for the data construction flow of LEX-ICON. (1) We manually collect 8,052 mimetic words and definitions from dictionaries, and systematically construct 2,930 disyllabic pseudo-words. (2) Using four LLMs (GPT-4.1, Qwen3-32B, Gemma-3-27B, and Gemini-2.5-flash), we automatically annotate each word with semantic dimensions based on its definitions. (3) For natural words, we retain features agreed upon by all models. For constructed words, we filter out features that are close to neutral. (4) The final dataset contains 10,982 words with 84,932 semantic features with varied input types.
We build LEX-ICON, a dataset with natural word group consisting of existing large-scale mimetic words derived from four natural languages (English, French, Japanese, and Korean), as well as constructed word group comprising systematically generated pseudo-words. Table 1 and 2 summarize the distribution of the overall datasets. Figure 2 also illustrates an overall dataset construction flow.
En. Fr. Ja. Ko. Con. Total
# Words 826 809 1418 4999 2930 10982
Table 1: Data distribution of LEX-ICON across natural (8,052 words from English, French, Japanese, and Korean) and constructed (2,930 words) groups. Japanese and Korean are known for their rich mimetic words (hamano1986sound; kwon2018iconicity), which contributes to their high proportion.
Dimension Natural Constructed Total
good-bad 2083 2083
beautiful-ugly 929 462 1391
pleasant-unpleasant 3380 3380
strong-weak 4299 208 4507
big-small 2073 1687 3760
rugged-delicate 2664 2664
active-passive 3884 3884
fast-slow 2051 1437 3488
sharp-round 2323 1623 3946
realistic-fantastical 6883 501 7384
structured-disorganized 3712 3712
ordinary-unique 1585 208 1793
interesting-uninteresting 208 208
simple-complex 4322 1602 5924
abrupt-continuous 4703 1005 5708
exciting-calming 2256 1402 3658
hard-soft 3676 1136 4812
happy-sad 719 1463 2182
harsh-mellow 3106 1005 4111
heavy-light 2918 1341 4259
inhibited-free 2673 206 2879
masculine-feminine 378 1522 1900
solid-nonsolid 2255 893 3148
tense-relaxed 2956 206 3162
dangerous-safe 448 541 989
Total 66484 18448 84932
Table 2: Semantic dimension distribution of pseudo ground truth data by word group. We adopt 25 semantic dimension criteria by sidhu2022higher to annotate LEX-ICON. Six dimensions from the constructed word group are excluded by removing close-to-neutral data points. As a result, 19 dimensions remain for experiments in §4.
3.1 Semantic Dimension
We employ the semantic dimension methodology, a concept originated from ``semantic differential'' by osgood1957measurement, which consists of two semantic feature adjectives located at opposite extremes, such as ``big'' and ``small.'' By applying this methodology to sound symbolism experiments, we can simultaneously and precisely measure the multifactorial meanings contained in a single word. We annotate up to 25 pairs of predefined semantic features as per sidhu2022higher for each word in the LEX-ICON, which is the most diverse scale in terms of studies on LLMs' iconicity.
3.2 Natural Mimetic Words
Data Collection.
We manually collect 8,052 mimetic words and definition data consisting of onomatopoeias and ideophones in English, French, Japanese, and Korean from specialized mimetic word lexicons and authoritative dictionaries for each language. We then extract the most representative definitions of these mimetic words from dictionaries such as the Oxford English Dictionary (dictionary1989oxford), Le Petit Robert (lepetitrobert2023), Nihon Kokugo Daijiten (9784095210216), and the Standard Korean Language Dictionary (nkl2025). For further source information, see the Appendix.
Input Type Variation.
To observe the effect of trained token memorization and modality change, we create three types of word form that contain textual and auditory input: (1) original text, (2) IPA-converted text with phoneme-level spacing, (3) audio waveform converted using text-to-speech (TTS) software. We perform IPA conversion with the Epitran package (Mortensen-et-al:2018), and obtain the TTS dataset using Google Text-to-Speech (google_tts_api_2025) for English, French, and Japanese; and MeloTTS (zhao2024melo) for Korean.
Large-Scale Annotation Process.
To effectively generate a large-scale ground truth data, all natural language words and their definitions in the LEX-ICON are given four LLMs: GPT-4.1 (openai2025_gpt41), Qwen3-32B (yang2025qwen3technicalreport) [^2] Non-reasoning mode., Gemma-3-27B (gemmateam2025gemma3technicalreport), and Gemini-2.5-flash (comanici2025gemini25pushingfrontier). The models are prompted to annotate each semantic dimension of a word with one of a feature pair or neutral labels (e.g., selecting one of the ``exciting'', ``calming'', or ``neither'' option) for each word. See the Appendix for the detailed prompt.
We finalize the results as pseudo ground truth by selecting unanimously agreed features across all models, deleting ``neither'' labels to remove the meaningless features for each word. As shown in Table 2, this process filters 67.0% of the total annotation points, yielding 66,484 high-quality semantic features for natural words. In §4, we verify these pseudo ground truth data through a human evaluation experiment. For more information on the semantic dimension ground truth, refer to the Appendix.
3.3 Constructed Mimetic Words
Phoneme Combination Generation.
We systematically construct novel words using a CVCV structure to create pseudo-words unlikely to be encountered during model training. We use 15 consonants from five categories: sonorants (/l/, /m/, /n/), voiced fricatives (/v/, /D/, /z/), voiceless fricatives (/f/, /s/, /S/), voiceless stops (/p/, /t/, /k/), and voiced stops (/b/, /d/, /g/). We use four vowels from two categories: front vowels (/i/, /ej/) and back vowels (/A/, /ow/). After filtering against existing entries in the IPA-dict database (ipa_dict) across four languages (US and UK English, French, Japanese, and Korean), we obtain 3,108 novel pseudo-words. We convert IPA symbols to English alphabet combinations for TTS compatibility and remove incorrectly pronounced words by Google TTS, yielding 2,930 final words.
Semantic Dimension Annotation.
We utilize the empirical coefficients from sidhu2022higher, who quantified associations between phoneme categories and semantic dimensions through human rating experiments. These coefficients represent how much phoneme categories deviate from the overall mean rating across 25 semantic dimensions. We assign corresponding scores to phonemes in our constructed words and calculate the mean score across the four phonemes. To focus on meaningful associations, we apply a threshold of 1.0 standard deviation from the neutral point, treating data points within this range and as semantically neutral, consistent with 'neither' labels used for multilingual mimetic words.
4 Semantic Dimension Prediction
We perform semantic dimension A/B tests to answer RQ 1, interpreting the results by semantic dimension, word group, and input type to quantitatively explore MLLMs' phonetic intuitions. Human evaluation is also conducted to ensure reliability of pseudo ground truth data in LEX-ICON, and the results are further compared to those of MLLM experiments. Details of the experiments are provided in the Appendix.
Figure 3: Macro-F1 score results for the semantic dimension A/B test. ``Natural'' and ``Constructed'' are results of LLM experiments, calculated by averaging all three input types (original text, IPA, and audio). Each dot represents each model's score for a given dimension. Human evaluation results only contain the ``Audio'' input type with sampled data for experimental feasibility, yet achieving superior scores compared to the baseline that demonstrate LEX-ICON's reliability.
4.1 Experimental Settings
We employ MLLMs that officially support simultaneous text and audio inputs: Qwen2.5-Omni (xu2025qwen25omnitechnicalreport) [^3] Qwen2.5-Omni-3B and Qwen2.5-Omni-7B., Gemini-2.5-flash (geminiteam2025geminifamilyhighlycapable), and GPT-4o (openai2024gpt4ocard)444gpt-4o for text-only input, gpt-4o-audio-preview for audio-enabled input.. All experimental models cover the four languages and orthographies present in LEX-ICON. For the Qwen models, inference is performed on one RTX 4090 GPU. We apply a zero-shot prompting strategy with temperature set to 0 throughout all experiments to ensure reproducibility.
4.2 Methodology
For each word, we prompt MLLMs with binary questions on each of the semantic dimensions and measure macro-F1 scores for the results of each dimension. We provide the models with words in three input types, keeping the query part as text. We insert audio tokens at the given words' positions within the prompt so that the audio tokens have the same series of positional embeddings with the text tokens, ensuring that the models infer internal representations in an integrated embedding space. An example prompt is illustrated in Table 3. The exact word input types are as follows:
  • Original text tokens (e.g., ``boom'').
  • Phoneme-level spaced IPA text tokens (e.g., ``b u m'').
  • TTS audio tokens (e.g., <AUDIO>555<AUDIO> represents a series of audio tokens that correspond to a given word.).
Semantic Dimension Test
Given an IPA [WORD] with its pronunciation audio, which semantic feature best describes the word based on auditory impression?
[WORD]
b u m (AUDIO: <AUDIO>)
[SEMANTIC DIMENSION]
exciting vs. calming
[OPTIONS]
1: exciting
2: calming
Answer with the number only. (1-2)
Table 3: An example of prompts for the semantic dimension A/B test in §4.2. This example illustrates a case of a natural language group English word ``boom'' that both the IPA text tokens and audio tokens (<AUDIO>) are presented. Detailed prompts are provided in the Appendix.
We calculate macro-F1 scores to mitigate the imbalance in simple accuracy for each model across all combinations of word groups, input types, and semantic dimensions. Results for ``natural'' group are averaged equally across the four languages. Refer to the Appendix for detailed methods.
4.3 Result
Phonetic Intuition by Semantic Dimension.
In Figure 3, MLLMs' macro-F1 scores averaged across all input types surpass the baseline score (0.50) in 84.2% (natural group) and 68.4% (constructed group) of the semantic dimensions, indicating that the models can detect phonetic iconicity not only in natural mimetic words that may have been memorized during training phase, but also in constructed pseudo-words with maximized sound-symbolic effects. Overall performance becomes even larger when the comparatively small-scaled Qwen2.5-Omni-3B model is excluded. These results are supported by human evaluation results that score above the baseline in most dimensions, which guarantees the reliability of our pseudo ground truth data automatically annotated by LLMs from dictionary data in §3.1. Notably, the models' strong performance on the sharp vs. round dimension aligns with well-known cognitive linguistic experiments like the ``bouba-kiki'' effect (ramachandran2001synaesthesia).
Human-like Iconicity.
Figure 4 shows that the Qwen2.5-Omni-7B model achieves the highest overall Pearson correlation coefficient with human evaluations across semantic dimensions, whereas larger models such as gemini-2.5-flash deviate more from human results. This suggests that while MLLMs can partially capture phonetic iconicity, they are still far from human-like semantic alignment. In particular, the relatively low correlation of IPA-converted natural word results indicates that linguistic arbitrariness from diverse languages may override iconic patterns even in mimetic words, as models' knowledge has been shaped by large-scale distributional semantics of natural languages, possibly leading to a tendency to overlook subtle phonosemantic cues.
Figure 4: Pearson correlation scores with human evaluation results by word group and input type. Higher scores reflect greater similarity to humans' semantic dimension score distributions, where Qwen2.5-Omni-7B scores the highest correlation (maximum $r=0.579$). In all models, constructed words elicit responses that are closer to human tendencies than natural words.
Linguistic Implication.
Our findings reveal systematic modality preferences across semantic dimensions, providing computational evidence for the multi-mechanism nature of sound symbolism, as shown in Figure 5. For dimensions where acoustic features are theoretically central, such as size distinctions (big vs. small) that correlate with formant frequencies (knoeferle2017drives) and speed distinctions (fast vs. slow) that relate to consonant voicing duration (saji2013cross), the MLLMs show a pattern of enhanced performance when processing constructed words in audio format, consistent with the theoretical predictions. Conversely, for dimensions where non-acoustic mechanisms are proposed, such as shape associations (sharp vs. round) based on lip rounding gestures (imai2025does) and valence associations (beautiful vs. ugly, happy vs. sad) affected by articulatory properties (korner2022articulation), the models exhibit stronger reliance on textural representations.
Figure 5: Advantage scores (macro-F1 differences) of audio inputs over the original text inputs by word group. X-axis indicates audio advantage scores for natural words, while Y-axis stands for constructed words. Each dot represents one semantic dimension, reflecting patterns aligned with linguistic implications, with an overall correlation (Pearson $r=0.681$, Spearman $\rho=0.705$).
5 Internal Attention Analysis
To address RQ 2, internal attention analysis focuses on the internal phenomena that emerge during the MLLM inference process. Then these phenomena are compared with linguistic theories to evaluate their correspondence.
Figure 6: Attention fraction scores indicating the ratio of semantic dimensions that the model focuses on for each IPA, by word group and input type. For each IPA and semantic dimension, the model tends to attend more to semantic features that exhibit stronger associations with the given phonemes in linguistics. For instance, the model mostly associates sharp semantic feature with /p/ and /k/, round with /m/ and /n/, big with /A/, and small with /i/ (kohler1967gestalt; parise2012audiovisual).
5.1 Experimental settings
We utilize the Qwen2.5-Omni-7B model for the experiment due to its performance correlation most similar to humans, as well as its accessibility of model weights, which enables direct analysis of internal representations. All other settings for the experiment are as in §4 to maintain consistency.
5.2 Methodology
We calculate which semantic feature the model pays more attention to, for each IPA symbol given a semantic dimension. We perform layer-wise analyses of relative attention scores for each IPA symbol across contrasting semantic features, distinguishing by input types (IPA text and audio waveform) and word groups (natural and constructed words). Unlike original text tokens, the IPA text and audio input types facilitate phoneme-level investigation, as each phoneme is represented by at least one token.
Attention Fraction Score Extraction.
For each inference conducted on the binary questions described in §4, we obtain the attention scores only when the model generates the correct response. For each layer, we calculate attention scores between tokens corresponding to a single IPA symbol (e.g., /w/ from /w I z/) and tokens corresponding to each semantic feature (e.g., fast vs. slow). After retrieving scores, we normalize the paired scores so that the sum of the two semantic feature scores is equal to one. These fraction scores undergo head-wise and word-wise averaging within each layer to yield a single mean score per layer. To mitigate the bias of assigning higher attention scores to preceding tokens, we repeat each experiment with the semantic features presented in reversed order and compute the mean normalized attention fraction score across both feature-order conditions.
Token-IPA Symbol Alignment.
When an IPA symbol spans multiple text tokens, the procedure first sums the attention scores across those tokens before fraction normalization. For audio inputs, we employ Montreal Forced Aligner (mcauliffe2025montreal) to segment each audio waveform into phonemes over time. The phoneme sequence is then aligned with the model’s 40 ms sampling period, with consecutive occurrences of the same phoneme treated as a single IPA symbol. Further details are in the Appendix.
5.3 Result
An attention fraction score above 0.5 for a given IPA-semantic feature pair indicates the model's preferential focus on phonemes with sound-symbolic associations. For more details about the experiment, refer to the Appendix.
Layer-wise Attention Fraction Score.
Figure 7 demonstrates that, for constructed words, the attention fraction scores for IPA text consistently exceed those for audio input type across layers (IPA = 0.523, audio = 0.506 on average), showing an upward trend toward the late layers. These lower phoneme‐level attention scores on audio inputs may imply that multimodal models derive greater benefits from extensively trained texts than from the acoustic properties of less-trained audio data. On the other hand, the average attention fraction scores for natural words (IPA = 0.507, audio = 0.501 on average) are lower than those for constructed words. This phenomenon may occur because arbitrary form–meaning mappings of natural words attenuate phonosemantic cues, thereby obscuring iconic phonemes. This interpretation is further corroborated by Figure 4, which shows the low correlation between human evaluation scores for natural words in both IPA text and audio modalities.
Phoneme-Semantic Feature Relation.
Figure 6 presents heatmaps of attention fraction scores for canonical IPA symbols and semantic dimensions by input type and word group. The patterns for constructed words in the IPA modality closely mirror prior findings in sound symbolism research. For example, phonemes such as /p/ and /k/ exhibit elevated attention under the sharp feature, whereas /m/ and /n/ associate with the round feature (kohler1967gestalt).
Figure 7: Attention fraction scores for the constructed word group, averaged across semantic dimensions. X-axis indicates the model's attention layer number, where Y-axis means the ratio at which the model attend to the correct semantic feature for a phoneme. While most layers score above the baseline fraction (0.50), the model focus on IPA text more than audio input for given iconic phonemes. Furthermore, the model tend to concentrate more on iconic phonemes in its late-layers.
6 Conclusion
In this work, we investigate MLLMs' phonetic iconicity on natural and constructed words via semantic dimension and internal attention analysis, constructing LEX-ICON, a large-scale mimetic word dataset for MLLM analysis for the first time. We discover that MLLMs have the ability to detect sound symbolism in both natural and constructed mimetic words, and pay a higher rate of attention in the internal layers to iconic phonemes. These results suggest that the models' sound-meaning association can be explained in terms of interpretability. Future work could further deepen the analytical methodology presented in this work through experiments with more human participants, investigate modality-specific information such as intonation, or extend it to application fields such as language learning or brand effects.
Acknowledgments
This work was partly supported by an Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2021-II211343, Artificial Intelligence Graduate School Program (Seoul National University), No.RS-2025-02263598, Development of Self-Evolving Embodied AGI Platform Technology through Real-World Experience), the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(RS-2024-00354218, RS-2024-00353125). We express special thanks to KAIT GPU project. The ICT at Seoul National University provides research facilities for this study.
Appendix A Details on LEX-ICON
A.1 Mimetic Word Data
Natural Word Collection.
We manually extract headwords and definitions from published onomatopoeia dictionaries for each language to determine whether a word belongs to the mimetic word category. For Korean, the word definition guideline of the NIKL_StandardKoreanLanguageDictionaryCompilationGuideline2_2000 is used to extract mimetic words and definitions. Semantic definitions that need to be supplemented are obtained from authoritative dictionaries for each language. The full list of the dictionaries is provided in Table 4.
Language References
English Rhyme over Reason: Phonological Motivation in English (benczes2019rhyme)
Oxford English Dictionary, 2nd Edition (dictionary1989oxford)
French Dictionnaire des Onomatopées (enckell2003dictionnaire)
Le Petit Robert de la Langue Française (lepetitrobert2023)
Trésor de la Langue Française (letresor1994)
Japanese Kurashi no Kotoba Gion Gitaigo Jiten (1130282271732203008)
Seisenban Nihon Kokugo Daijiten (9784095210216)
Kojien, 7th Edition (kojien7)
Digital Daijisen (digitaldaijisen)
Daijirin, 4th Edition (matsumura2019daijirin)
Korean Standard Korean Language Dictionary (nkl2025)
Table 4: Reference materials for natural mimetic word collection by language.
We filter the data obtained from the mimetic word dictionaries, consisting of 954 English words, 1049 French words, 2025 Japanese words, and 4999 Korean words, using information from the aforementioned authoritative dictionaries. We finally determine 826 English words, 809 French words, 1418 Japanese words, and 4999 Korean words as natural mimetic word data.
Constructed Word Generation.
We apply systematic pre-processing to IPA-dict entries from open-dict-data dictionaries during the filtering process to exclude words that have the same pronunciation as natural language words. We remove non-distinctive markers including stress (”) and duration (:) symbols, and convert allophones to their corresponding phonemes such as changing velarized lateral [] to /l/ in English and tap [R] to /l/ in Korean. We also standardize notation by converting /eI/ to /ej/, /oU/ to /ow/.
For TTS compatibility, we create IPA-to-alphabet mapping rules and refine mappings where needed, which includes converting ``tho'' to ``though'' to ensure correct /Dow/ pronunciation. Generated words are hyphenated (e.g., ``lah-mo'') except for combinations ending in /Di/ (e.g., ``laythey'').
Word Group.
In Figure 8, we illustrate example words of the natural and constructed word groups that comprise LEX-ICON.
Figure 8: An illustrative description of the word groups. Natural word group consists of English, French, Japanese, and Korean mimetic words. Constructed word group includes disyllabic pseudo-words that experimental models may not have trained.
Input Type.
We convert each word into three types (original orthographic text, phoneme-level spaced IPA text, and text-to-speech audio waveform) and input them into the experimental models. Figure 9 shows the semantic dimension prediction performance of models for each word group and input type.
Figure 9: Semantic dimension prediction performance of models. For each input type, the dots for natural words' results are on the left, while constructed words' results are on the right. Each dot indicates macro-F1 scores of the models' response averaged across semantic dimensions.
A.2 Semantic Dimension
Table 6 shows the extreme feature pairs of 25 semantic dimensions suggested by sidhu2022higher, which have been utilized for our semantic dimension prediction experiments.
After the automatic annotation process for natural words is complete, we utilize labels that are agreed upon by all model families and not annotated as ``neither'' as pseudo ground truth. Similarly, we remove semantic features close to neutral from the dataset by deleting semantic dimensions from each constructed word whose semantic features' ground truth score is within 1.0 standard deviation of the overall distribution. Due to this method, words in the datasets have fewer than 25 dimension labels. The distribution of these labels is shown in Table 5.
Dimension Natural Constructed Total
(Feature 1)-(Feature 2) Ft. 1 Ft. 2 Total Ft. 1 Ft. 2 Total Ft. 1 Ft. 2 Total
good-bad 532 1551 2083 0 0 0 532 1551 2083
beautiful-ugly 290 639 929 344 118 462 634 757 1391
pleasant-unpleasant 797 2583 3380 0 0 0 797 2583 3380
strong-weak 2273 2026 4299 107 101 208 2380 2127 4507
big-small 902 1171 2073 894 793 1687 1796 1964 3760
rugged-delicate 1327 1337 2664 0 0 0 1327 1337 2664
active-passive 3267 617 3884 0 0 0 3267 617 3884
fast-slow 1579 472 2051 750 687 1437 2329 1159 3488
sharp-round 1644 679 2323 789 834 1623 2433 1513 3946
realistic-fantastical 6876 7 6883 251 250 501 7127 257 7384
structured-disorganized 705 3007 3712 0 0 0 705 3007 3712
ordinary-unique 229 1356 1585 101 107 208 330 1463 1793
interesting-uninteresting 135 73 208 0 0 0 135 73 208
simple-complex 2064 2258 4322 784 818 1602 2848 3076 5924
abrupt-continuous 2055 2648 4703 440 565 1005 2495 3213 5708
exciting-calming 1741 515 2256 693 709 1402 2434 1224 3658
hard-soft 1757 1919 3676 440 696 1136 2197 2615 4812
happy-sad 436 283 719 709 754 1463 1145 1037 2182
harsh-mellow 2477 629 3106 440 565 1005 2917 1194 4111
heavy-light 1048 1870 2918 622 719 1341 1670 2589 4259
inhibited-free 1484 1189 2673 105 101 206 1589 1290 2879
masculine-feminine 144 234 378 689 833 1522 833 1067 1900
solid-nonsolid 933 1322 2255 105 788 893 1038 2110 3148
tense-relaxed 1986 970 2956 105 101 206 2091 1071 3162
dangerous-safe 392 56 448 440 101 541 832 157 989
Table 5: Semantic dimension distribution of pseudo ground truth by word group.
Feature A Feature B
good bad
beautiful ugly
pleasant unpleasant
strong weak
big small
rugged delicate
active passive
fast slow
sharp round
realistic fantastical
structured disorganized
ordinary unique
interesting uninteresting
simple complex
abrupt continuous
exciting calming
hard soft
happy sad
harsh mellow
heavy light
inhibited free
masculine feminine
solid nonsolid
tense relaxed
dangerous safe
Table 6: The 25 semantic dimension list by sidhu2022higher.
A.3 An Example Entry from LEX-ICON
LEX-ICON consists of the following information for each mimetic word, as shown in Table 7.
Key Value
word whizz
meaning To make a sound as of a body rushing through the air.
ref Rhyme Over Reason: Phonological Motivation in English
ipa w I z
romanization whizz
language en
(a) Metadata from an example word.
Dimension Feature
active-passive active
fast-slow fast
sharp-round sharp
realistic-fantastical realistic
simple-complex simple
abrupt-continuous continuous
exciting-calming exciting
harsh-mellow harsh
inhibited-free free
solid-nonsolid nonsolid
(b) Semantic dimension data from an example word.
Table 7: An example entry from LEX-ICON.
Appendix B Experimental Settings
B.1 Inference Settings
  • Hardware: four NVIDIA RTX 4090 GPUs for 27B and 32B models, one NVIDIA RTX 4090 GPU for 3B and 7B models, and API calls for other proprietary models.
  • Software: PyTorch (paszke2017automatic) and vLLM (kwon2023efficient) libraries.
  • Random seed: 42 (for reproducibility, only if needed).
  • Temperature: 0 (for reproducibility).
  • Maximum output tokens: 1024.
B.2 Language Models Used in This Work
In this work, we utilize the following language models for data annotation, experimentation, and analysis.
  • gpt-4o (openai2024gpt4ocard)
  • gpt-4o-audio-preview (openai2024gpt4ocard)
  • gpt-4.1 (openai2025_gpt41)
  • gemini-2.5-flash (comanici2025gemini25pushingfrontier)
  • gemma-3-27b-it (gemmateam2025gemma3technicalreport)
  • Qwen3-32B (yang2025qwen3technicalreport)
  • Qwen2.5-Omni-3B (xu2025qwen25omnitechnicalreport)
  • Qwen2.5-Omni-7B (xu2025qwen25omnitechnicalreport)
Appendix C Token-IPA Symbol Alignment
To enable phoneme-level attention analysis, we implement a pipeline for aligning audio tokens with IPA symbols.
C.1 Montreal Forced Alignment (MFA) Pipeline
We employ Montreal Forced Alignment (mcauliffe2025montreal) to obtain precise phoneme-level segmentation of audio waveforms. We utilize pre-trained acoustic models for each language:
  • English (US) ARPA acoustic model v3.0.0 (mfa_english_us_arpa_acoustic_2024)
  • French MFA acoustic model v3.0.0 (mfa_french_mfa_acoustic_2024)
  • Japanese MFA acoustic model v3.0.0 (mfa_japanese_mfa_acoustic_2024)
  • Korean MFA acoustic model v3.0.0 (mfa_korean_mfa_acoustic_2024)
The alignment process generates Textgrid files containing phoneme boundaries with millisecond precision.
C.2 Custom Dictionary Generation and OOV Handling
We generate language-specific pronunciation dictionaries by mapping word-IPA pairs from our dataset to MFA phoneme representations. When initial alignment fails due to out-of-vocabulary (OOV) words, we automatically (1) identify words producing `spn' (spoken noise) tokens in TextGrid output, (2) generate pronunciations using Epitran for OOV words, (3) update pronunciation dictionaries with new entries, and (4) re-run MFA alignment until convergence.
C.3 TextGrid Parsing and Audio Token Alignment
We parse MFA-generated TextGrid files to extract phoneme boundaries and convert them into discrete frame sequences compatible with the model's 40ms sampling period. For each audio frame, we assign the phoneme whose temporal boundary encompasses the frame's center point. The resulting frame-level sequence is then aligned with the model's audio token representations, enabling precise measurement of phoneme-level attention patterns for audio inputs.
Appendix D Metrics
D.1 Semantic Dimension Prediction
Semantic Dimension Macro-F1 Score.
  • $S$: Set of semantic dimensions, $|S|=d$.
  • $C$: Set of semantic features in each semantic dimension. $|C|=2$.
  • $TP_{s,c}$: True–positive count for dimension $s$ and feature $c$.
  • $FP_{s,c}$: False–positive count for dimension $s$ and feature $c$.
  • $FN_{s,c}$: False–negative count for dimension $s$ and feature $c$.
The macro-F1 score for each semantic dimension $s$ is defined as follows:
$\mathrm{MacroF1}_{s}\;=\;\frac{1}{|C|}\sum_{c\in C}\frac{2\,TP_{s,c}}{2\,TP_{s,c}+FP_{s,c}+FN_{s,c}}$
We calculate the final natural group scores for each semantic dimension of LLMs by averaging the macro-F1 scores for four natural languages (English, French, Japanese, and Korean) and three input types (original text, IPA text, and audio waveform) for each dimension. For the constructed words, only averaging three input types is performed with the same procedure as in the calculation of natural words' scores.
Correlation Score with Human Evaluation.
  • $D_{\text{valid}}$: the set of semantic dimensions for which both human evaluation macro-F1 scores and a model's macro-F1 scores are available.
  • $h_{d}$: the human evaluation macro-F1 score for dimension $d$ and audio input type.
  • $m_{d}$: the model's macro-F1 score for dimension $d$ and a given input type.
  • $\mu_{H}$: the mean of the human scores across all dimensions in $D_{\text{valid}}$.
  • $\mu_{M}$: the mean of the model's scores across all dimensions in $D_{\text{valid}}$.
The Pearson correlation coefficient $r$ between the human scores ($H$) and the model's scores ($M$) for a given word group and input type is calculated as:
$r(H,M)=\frac{\sum_{d\in D_{\text{valid}}}(h_{d}-\mu_{H})(m_{d}-\mu_{M})}{\sqrt{\sum_{d\in D_{\text{valid}}}(h_{d}-\mu_{H})^{2}}\sqrt{\sum_{d\in D_{\text{valid}}}(m_{d}-\mu_{M})^{2}}}$ (1)
Audio Input Type Advantage Score.
Let $S(g,t,d)$ be the macro F1-score for a given word group $g\in\{\text{natural, constructed}\}$, input type $t\in\{\text{audio, original text}\}$, and semantic dimension $d$.
The Input Type Advantage Score, denoted as $A(g,d)$, for a specific word group $g$ and semantic dimension $d$ is defined as the difference between the scores of the two input types:
$A(g,d)=S(g,\text{audio},d)-S(g,\text{original text},d)$ (2)
A positive value of $A(g,d)$ indicates that the ``audio'' input type has an advantage over the ``original text'' input type for that specific dimension and word group. Conversely, a negative value indicates an advantage for the ``original text'' input type.
The score $S(g,t,d)$ is calculated by averaging the macro F1-scores across all models for the given parameters:
$S(g,t,d)=\frac{1}{|K|}\sum_{k\in K}s_{k,g,t,d}$ (3)
Where:
  • $s_{k,g,t,d}$: the macro F1-score of an individual model $k$.
  • $K$: the set of all models evaluated.
  • $|K|$: the total number of models.
D.2 Internal Attention Analysis
Attention Fraction Score Computation.
The Phoneme-Semantic Feature Attention Fraction Score is calculated by the algorithm described in Algorithm 1.
Algorithm 1 Phoneme-Semantic Feature Attention Fraction Score Calculation.
1:Load model and prompt
[…]
6:
7:Data format: {wordword: [($feature_{1}$, $feature_{2}$), answeranswer]} $\triangleright$ feature_i: features from semantic dims
8:
9:Model Inference Phase
10:for all wordword, ($feature_{1}$, $feature_{2}$, answeranswer) in data do
11:  $ipa\_text\leftarrow$ to_ipa(wordword)
12:  $audio\leftarrow$ to_audio(wordword)
13:  for all $input\in\{ipa\_text,audio\}$ do
14:   $prompt\leftarrow$ format_prompt(inputinput, $feature_{1}$, $feature_{2}$) $\triangleright$ Format prompt using input and semantic features
15:   responseresponse, attnattn, $tokens\leftarrow$ inference(modelmodel, promptprompt) $\triangleright$ attn = [layer, head, query, key]
16:   idx_inputidx\_input, $idx\_feature_{1}$, $idx\_feature_{2}\leftarrow$ find_idx(tokenstokens, inputinput, $feature_{1}$, $feature_{2}$) $\triangleright$ type: list[int]
17:   $idxs\leftarrow\{idx\_input,idx\_feature_{1},idx\_feature_{2}\}$ $\triangleright$ Aggregate relevant token indices
18:   $filtered\_attn\leftarrow attn[:,:,idxs,idxs]$ $\triangleright$ Select sub-attention map for relevant tokens only
19:   save(filtered_attnfiltered\_attn, tokenstokens, inputinput, idx_inputidx\_input, $idx\_feature_{1}$, $idx\_feature_{2}$, responseresponse)
20:  end for
21:end for
22:
23:Phoneme-Semantic Feature Computation Phase
24:for all $input\in\{ipa\_text,audio\}$ do
25:  $result\leftarrow\{\}$
26:  for all wordword, ($feature_{1}$, $feature_{2}$, answeranswer) in data do
27:   filtered_attnfiltered\_attn, tokenstokens, inputinput, idx_inputidx\_input, $idx\_feature_{1}$, $idx\_feature_{2}$, $response\leftarrow$ load(wordword, $feature_{1}$, $feature_{2}$)
28:   if $answer\neq response$ then
29:     continue
30:   end if
31:   $ipa\_span\leftarrow$ span_aligning_input_to_tokens(inputinput, tokenstokens, idx_inputidx\_input)
32:   for all $layers\in filtered\_attn$ do
33:     for all $layer\in layers$ do $\triangleright$ layer = [head, query, key]
34:      for all $idx\_phoneme\in ipa\_span$ do
35:        $score_{1}\leftarrow\text{sum}(layer[:,idx\_feature_{1},idx\_phoneme])$
36:        $score_{2}\leftarrow\text{sum}(layer[:,idx\_feature_{2},idx\_phoneme])$
37:        $feature_{1score}$ $\leftarrow$ $score_{1}/(score_{1}+score_{2})$
38:        $feature_{2score}$ $\leftarrow$ $score_{2}/(score_{1}+score_{2})$
39:        if $result[phoneme][feature_{1}][layer]$ is undefined then
40:         $result[phoneme][feature_{1}][layer]\leftarrow[\,]$
41:        end if
42:        if $result[phoneme][feature_{2}][layer]$ is undefined then
43:         $result[phoneme][feature_{2}][layer]\leftarrow[\,]$
44:        end if
45:        append($result[phoneme][feature_{1}][layer]$, $feature_{1score}$)
46:        append($result[phoneme][feature_{2}][layer]$, $feature_{2score}$)
47:      end for
48:     end for
49:   end for
50:  end for
51:  for all $phoneme\in ipa\_list$ do
52:   for all $feature\in dims$ do
53:     for all $layer\in result[phoneme][feature]$ do
54:      $result[phoneme][feature][layer]\leftarrow\text{mean}(result[phoneme][feature][layer])$
55:     end for
56:   end for
57:  end for
58:  save(result)
59:end for
Appendix E Human Evaluation
To ensure the reliability of the semantic dimension selection and the automatic annotation of the pseudo ground truth data in LEX-ICON, we conduct a human evaluation using 152 questions randomly sampled for each participant from the entire dataset. 10 graduate students without expertise in linguistics participate in this experiment, which is performed via the Label Studio (LabelStudio). The experiment utilize questions with the same prompt structure as the semantic dimension prediction test that was provided to LLMs. With regard to the experiment's feasibility, all words are presented in audio waveforms directly hearable to humans. An example image of the process is displayed in Figure 10.
The randomly sampled test data contains 19 semantic dimensions that are presented in both of the natural and constructed word group data. For each dimension, all natural language types and answer semantic feature distributions are selected from the same distribution, respectively. As a result, four natural group words and four constructed group words are extracted for each dimension.
Due to the nature of sound symbolism task, in which disagreements may arise between responses, the human response distribution may have a wider variance than typical reasoning-centric experiments. However, we maximize reliability within limited resources through the aforementioned methodology. Table 8 shows human response average macro-F1 scores and standard deviation of the scores by semantic dimension.
Figure 10: Screenshot image of the human evaluation process. Human evaluators are presented with the same prompts provided to LLMs via the Label Studio interface. Participants select a semantic feature after listening to audio waveform.
Dimension Macro-F1 (± Std.)
tense-relaxed 0.6784(± 0.2052)0.6784\,($\pm\,$0.2052)
realistic-fantastical 0.6333(± 0.2080)0.6333\,($\pm\,$0.2080)
masculine-feminine 0.6287(± 0.1643)0.6287\,($\pm\,$0.1643)
fast-slow 0.6210(± 0.2185)0.6210\,($\pm\,$0.2185)
big-small 0.6058(± 0.1298)0.6058\,($\pm\,$0.1298)
happy-sad 0.6003(± 0.1303)0.6003\,($\pm\,$0.1303)
abrupt-continuous 0.5903(± 0.2205)0.5903\,($\pm\,$0.2205)
beautiful-ugly 0.5844(± 0.1493)0.5844\,($\pm\,$0.1493)
heavy-light 0.5808(± 0.1648)0.5808\,($\pm\,$0.1648)
exciting-calming 0.5797(± 0.1948)0.5797\,($\pm\,$0.1948)
sharp-round 0.5796(± 0.1072)0.5796\,($\pm\,$0.1072)
hard-soft 0.5794(± 0.2142)0.5794\,($\pm\,$0.2142)
solid-nonsolid 0.5682(± 0.2259)0.5682\,($\pm\,$0.2259)
inhibited-free 0.5621(± 0.2263)0.5621\,($\pm\,$0.2263)
dangerous-safe 0.5594(± 0.1887)0.5594\,($\pm\,$0.1887)
harsh-mellow 0.5536(± 0.1652)0.5536\,($\pm\,$0.1652)
ordinary-unique 0.5376(± 0.1891)0.5376\,($\pm\,$0.1891)
strong-weak 0.4863(± 0.2000)0.4863\,($\pm\,$0.2000)
simple-complex 0.4602(± 0.1752)0.4602\,($\pm\,$0.1752)
Table 8: Human evaluation result with ranked average macro-F1 score and standard deviation by semantic dimension. Among the 25 semantic dimensions, 19 were used for human evaluation due to their presence in both natural and constructed word groups and their sufficient frequency in the dataset.
Appendix F Detailed Results
F.1 Semantic Dimension Prediction Experiment
Table 9, 10, 11, and 12 show accuracies and macro-F1 scores of experimental models' results for each input types and semantic dimensions. In the tables, ``–'' denotes a dimension where all ground truth features are classified as ``neither'' thus removed.
Dimension Natural Constructed
Original IPA Audio Original IPA Audio
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
good-bad 67.0 61.9 38.9 38.4 53.1 48.8
beautiful-ugly 68.6 62.6 28.6 25.5 48.9 43.1 72.9 60.5 77.1 54.3 68.0 50.6
pleasant-unpleasant 68.4 63.6 43.9 43.0 56.5 53.1
strong-weak 71.0 68.2 55.8 52.1 66.4 62.6 60.1 60.0 52.4 51.9 65.9 64.3
big-small 61.6 54.3 62.5 60.9 60.5 54.1 49.1 39.0 64.4 64.0 50.3 46.9
rugged-delicate 65.9 64.8 50.1 48.7 59.9 58.3
active-passive 72.4 54.3 62.8 45.8 71.0 51.9
fast-slow 66.7 56.9 54.6 46.0 61.0 53.4 58.5 58.5 64.4 62.1 67.7 67.6
sharp-round 76.6 72.1 62.7 59.2 69.8 65.4 53.7 53.7 73.1 73.0 58.4 58.0
realistic-fantastical 68.0 39.8 19.7 15.9 75.3 43.0 60.3 53.4 58.5 51.8 79.8 79.8
structured-disorganized 69.8 65.0 24.7 21.5 57.4 54.8
ordinary-unique 76.8 50.8 74.3 48.0 68.2 48.9 51.9 35.9 64.4 58.2 63.9 61.0
interesting-uninteresting 72.9 63.2 66.0 38.7 67.6 45.6
simple-complex 69.0 44.1 70.8 49.1 68.2 41.2 49.8 34.6 49.0 33.0 48.8 33.2
abrupt-continuous 75.1 71.5 58.5 54.8 66.9 64.7 46.8 42.8 61.0 54.8 53.9 53.6
exciting-calming 51.3 47.3 22.1 19.9 42.9 39.9 52.6 41.3 59.9 52.0 53.6 48.4
hard-soft 71.7 70.9 55.2 52.8 64.7 63.1 57.4 53.5 64.8 58.7 57.2 55.8
happy-sad 75.8 74.4 54.4 48.4 62.2 58.2 70.3 70.0 56.7 51.1 69.0 68.1
harsh-mellow 84.9 68.8 55.8 48.7 78.5 61.2 54.7 51.3 67.3 66.8 51.5 46.7
heavy-light 74.7 73.7 56.1 55.3 66.9 65.7 64.8 64.5 50.4 47.8 46.1 43.2
inhibited-free 69.1 67.3 54.0 53.4 63.7 61.2 32.0 31.8 30.6 30.2 41.7 35.9
masculine-feminine 65.6 61.8 39.3 36.9 57.9 55.8 62.2 62.2 66.6 66.3 68.0 68.0
solid-nonsolid 75.8 69.1 63.1 55.5 70.7 61.1 87.5 52.9 77.4 56.2 81.1 60.6
tense-relaxed 74.7 66.8 56.5 50.9 64.6 56.6 55.8 55.8 77.7 77.5 65.5 64.6
dangerous-safe 70.5 58.7 54.8 43.0 53.6 45.3 61.6 49.4 57.9 50.0 59.0 52.7
Table 9: Detailed semantic dimension prediction accuracy and macro-F1 score results for Qwen2.5-Omni-7B.
Dimension Natural Constructed
Original IPA Audio Original IPA Audio
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
good-bad 70.3 62.7 60.1 49.3 60.0 55.0
beautiful-ugly 72.6 64.6 64.0 47.3 60.9 56.1 43.7 40.5 38.7 38.7 67.5 52.1
pleasant-unpleasant 64.0 59.7 50.3 46.2 42.5 40.7
strong-weak 53.6 53.4 38.4 27.4 62.8 54.2 49.0 34.5 48.6 32.7 62.0 55.6
big-small 57.1 47.9 51.7 35.8 56.4 46.8 46.7 32.2 47.4 34.1 51.4 44.1
rugged-delicate 67.9 67.2 44.5 33.9 63.4 62.2
active-passive 49.9 42.8 14.4 12.5 76.2 53.0
fast-slow 51.1 47.2 26.1 23.8 53.2 47.0 58.9 55.6 49.1 39.6 61.7 61.7
sharp-round 77.2 69.4 63.3 59.4 80.3 70.4 55.5 53.8 73.2 72.7 54.3 47.2
realistic-fantastical 18.7 15.6 0.6 0.6 20.2 16.5 52.3 38.4 50.1 33.7 51.9 40.7
structured-disorganized 70.6 62.7 44.9 43.9 51.4 50.1
ordinary-unique 68.6 60.6 74.8 51.3 54.2 49.4 66.8 63.7 66.3 61.8 64.4 64.2
interesting-uninteresting 70.4 58.5 50.2 42.4 68.1 48.3
simple-complex 70.2 57.5 51.7 41.0 69.4 51.9 56.9 53.2 60.4 56.1 49.3 34.7
abrupt-continuous 69.2 66.8 46.2 34.7 63.7 60.2 42.7 34.0 55.9 35.9 54.7 48.8
exciting-calming 58.6 51.4 23.5 20.9 47.9 43.5 71.5 71.4 62.0 56.0 62.9 61.8
hard-soft 57.6 50.4 48.8 32.8 54.1 44.5 61.2 38.2 61.3 38.0 61.4 39.9
happy-sad 68.7 66.4 55.3 52.6 66.2 60.0 68.6 68.1 39.6 35.5 52.4 44.5
harsh-mellow 67.9 57.9 51.0 45.1 58.2 49.3 47.4 42.7 53.4 50.4 55.6 52.7
heavy-light 64.0 57.8 55.9 35.7 60.1 49.7 53.4 35.4 53.9 35.6 54.3 37.3
inhibited-free 61.6 54.4 53.0 34.4 56.9 43.7 47.6 33.0 49.0 32.9 48.1 32.5
masculine-feminine 71.0 64.0 44.2 32.2 66.9 60.2 58.0 49.8 56.9 41.3 71.5 70.7
solid-nonsolid 68.9 49.4 67.1 39.9 67.8 48.5 87.0 47.4 88.2 46.9 84.3 49.1
tense-relaxed 62.2 58.0 31.1 27.0 55.1 50.5 63.1 63.1 64.1 60.0 68.9 68.9
dangerous-safe 43.2 40.4 15.2 13.8 32.9 30.0 19.0 16.3 19.6 17.0 25.0 23.9
Table 10: Detailed semantic dimension prediction accuracy and macro-F1 score results for Qwen2.5-Omni-3B.
Dimension Natural Constructed
Original IPA Audio Original IPA Audio
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
good-bad 76.1 71.4 48.9 46.8 54.2 51.3
beautiful-ugly 82.7 78.0 50.3 46.3 56.7 52.4 74.5 47.2 77.3 53.9 72.1 46.0
pleasant-unpleasant 75.8 71.3 50.4 48.0 51.5 49.8
strong-weak 78.8 77.5 65.7 61.7 64.3 63.3 77.4 77.4 67.8 67.7 77.9 77.9
big-small 83.4 82.1 69.4 68.4 68.8 65.9 63.0 59.3 75.3 74.8 64.5 62.2
rugged-delicate 80.7 79.8 65.7 64.0 67.7 65.7
active-passive 86.3 63.7 77.9 50.2 69.3 52.3
fast-slow 89.1 75.7 78.1 58.3 65.2 55.1 66.9 62.0 71.1 68.5 72.2 72.0
sharp-round 85.4 81.8 75.2 70.3 59.0 56.5 70.9 69.7 77.2 76.9 62.5 57.5
realistic-fantastical 23.9 18.8 13.4 10.7 32.5 23.6 50.5 35.3 39.3 30.7 63.1 61.9
structured-disorganized 83.9 76.7 37.2 36.8 52.3 51.1
ordinary-unique 77.4 51.0 77.6 58.4 68.3 46.8 51.0 33.8 61.1 53.3 51.9 43.1
interesting-uninteresting 70.9 50.6 68.8 52.7 70.4 58.5
simple-complex 73.1 58.2 74.0 57.3 71.5 55.5 52.7 41.0 49.9 35.1 54.6 46.3
abrupt-continuous 80.5 74.6 67.3 62.8 71.5 67.6 60.3 57.9 64.8 64.2 67.5 67.3
exciting-calming 84.1 73.5 55.9 50.8 49.8 45.4 66.2 65.9 74.0 73.3 54.4 46.7
hard-soft 80.5 80.2 64.5 63.9 67.2 65.2 63.4 45.4 71.6 64.1 71.2 64.2
happy-sad 79.4 74.7 58.8 52.1 60.1 58.0 49.4 35.5 51.9 40.2 62.5 61.5
harsh-mellow 75.5 65.5 53.1 47.0 47.8 43.5 59.8 47.6 60.6 46.4 62.2 52.7
heavy-light 81.6 79.6 70.6 69.3 68.1 61.7 57.4 43.6 73.9 71.0 55.6 40.8
inhibited-free 73.6 71.4 64.4 61.9 60.0 52.4 50.5 38.1 51.9 40.9 47.1 33.5
masculine-feminine 83.5 77.4 68.3 58.4 71.2 63.4 69.4 67.0 78.1 78.0 82.7 82.2
solid-nonsolid 81.0 76.8 63.0 60.8 69.5 66.3 87.9 60.2 67.7 59.3 70.8 61.4
tense-relaxed 79.4 73.4 66.3 57.6 68.0 61.8 71.8 71.7 68.0 66.3 85.9 85.9
dangerous-safe 63.2 54.9 45.0 39.9 28.8 26.0 40.9 40.6 27.7 27.1 19.4 16.7
Table 11: Detailed semantic dimension prediction accuracy and macro-F1 score results for gpt-4o.
Dimension Natural Constructed
Original IPA Audio Original IPA Audio
Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1 Acc. F1
good-bad 86.0 80.2 76.4 57.7 78.0 53.5
beautiful-ugly 90.7 86.7 72.2 60.1 62.4 55.7 72.7 65.1 67.1 64.3 39.4 38.0
pleasant-unpleasant 82.2 76.8 71.4 59.7 79.2 52.3
strong-weak 76.7 73.8 60.5 58.9 62.9 56.3 62.0 61.5 66.8 66.0 47.6 45.7
big-small 83.3 82.8 65.0 61.0 59.6 55.2 63.0 59.4 62.5 59.0 50.5 50.0
rugged-delicate 83.2 82.3 68.0 65.9 66.0 56.7
active-passive 87.5 59.8 70.2 49.6 68.2 49.6
fast-slow 80.7 70.0 47.6 44.1 30.9 30.5 63.9 63.9 58.0 51.9 50.9 41.5
sharp-round 86.5 82.2 73.3 68.8 76.9 65.3 61.2 60.2 73.1 71.9 56.6 52.1
realistic-fantastical 69.6 40.9 33.5 24.7 86.4 46.4 51.1 48.4 67.3 66.3 55.7 53.1
structured-disorganized 84.2 78.2 67.0 58.0 69.4 56.8
ordinary-unique 80.7 72.3 68.6 56.3 51.3 46.4 54.8 46.0 62.0 61.5 56.7 55.9
interesting-uninteresting 74.2 59.9 62.7 47.7 62.3 51.0
simple-complex 80.8 68.4 67.8 57.2 54.6 47.9 57.9 51.6 58.7 57.0 49.9 49.9
abrupt-continuous 84.2 81.3 73.0 69.6 71.7 61.5 55.4 55.3 61.7 57.5 50.0 48.3
exciting-calming 84.5 71.8 65.6 54.1 58.2 50.7 54.8 50.1 71.8 71.7 56.2 55.7
hard-soft 79.8 79.6 64.2 62.7 59.6 58.3 63.6 58.2 67.3 56.6 53.1 51.9
happy-sad 81.3 80.3 57.2 55.7 51.0 42.2 59.7 59.5 60.7 58.8 53.0 39.2
harsh-mellow 85.6 73.8 70.1 57.1 75.5 56.1 59.9 57.6 60.7 51.9 49.5 46.7
heavy-light 82.2 81.7 70.0 69.4 56.4 48.8 66.8 61.3 75.3 73.3 47.1 35.0
inhibited-free 74.4 73.9 60.9 60.1 55.8 51.6 46.6 44.7 47.1 44.4 50.0 40.9
masculine-feminine 82.5 76.2 59.6 52.7 54.0 51.7 67.0 61.8 65.0 60.5 58.1 58.0
solid-nonsolid 84.8 82.0 68.0 60.8 47.5 46.2 88.2 61.1 87.9 73.4 25.6 25.6
tense-relaxed 76.3 71.4 64.0 55.8 69.4 58.5 61.7 61.0 79.6 79.5 65.5 62.2
dangerous-safe 78.8 65.3 73.7 52.3 80.3 52.2 66.9 56.4 53.6 51.3 77.3 46.6
Table 12: Detailed semantic dimension prediction accuracy and macro-F1 score results for gemini-2.5-flash.
F.2 Internal Attention Analysis
We provide full result tables for the Phoneme-Semantic Feature Attention Fraction Scores by word group (natural and constructed) and input type (IPA text and audio waveform).
  • Natural - IPA text: Table 13, 14, 15, and 16.
  • Natural - Audio waveform: Table 17, 18, 19, and 20.
  • Constructed - IPA text: Table 21 and 22.
  • Constructed - Audio waveform: Table 23 and 24.
Dimension i I e E æ a W u U o 2 O A @
abrupt .506 .491 .512 .511 .510 .517 .503 .506 .424 .509 .510 .492 .506 .511
continuous .494 .509 .488 .489 .490 .482 .497 .494 .576 .490 .490 .508 .493 .489
active .495 .499 .493 .482 .507 .502 .469 .496 .407 .496 .493 .461 .505 .492
passive .505 .501 .507 .517 .493 .498 .530 .504 .593 .504 .507 .539 .495 .508
beautiful .496 .455 .462 .499 .425 .487 .470 .443 .506 .484 .426 .442 .405 .486
ugly .504 .545 .538 .501 .575 .513 .529 .556 .494 .516 .573 .558 .594 .514
big .492 .473 .551 .452 .495 .469 .497 .515 .501 .465 .559 .531 .559 .506
small .507 .527 .449 .547 .505 .531 .503 .485 .499 .535 .441 .469 .440 .494
dangerous .507 .517 .512 .513 .528 .503 .505 .510 .507 .521 .505 .521 .510
safe .493 .483 .488 .486 .471 .496 .495 .490 .493 .479 .494 .478 .490
exciting .480 .484 .470 .478 .469 .471 .449 .474 .474 .468 .461 .442 .457
calming .520 .516 .529 .521 .531 .527 .550 .526 .526 .531 .538 .558 .543
fast .488 .570 .499 .521 .575 .519 .466 .478 .480 .492 .516 .534 .514
slow .512 .429 .501 .479 .425 .481 .533 .522 .519 .507 .483 .465 .486
good .500 .458 .468 .497 .417 .470 .487 .486 .591 .487 .475 .481 .453 .467
bad .500 .542 .532 .503 .583 .529 .513 .514 .409 .512 .525 .519 .547 .533
happy .522 .520 .484 .530 .471 .512 .508 .482 .469 .513 .517 .530 .502
sad .478 .479 .516 .470 .528 .488 .492 .517 .530 .486 .482 .469 .498
hard .479 .478 .476 .476 .484 .489 .473 .483 .478 .483 .470 .480 .475
soft .521 .522 .524 .524 .516 .510 .526 .517 .522 .517 .530 .520 .525
harsh .279 .266 .276 .262 .261 .283 .275 .283 .256 .273 .277 .275 .280 .258
mellow .721 .734 .724 .738 .739 .717 .725 .717 .744 .727 .723 .725 .720 .741
heavy .473 .459 .492 .466 .489 .471 .487 .500 .510 .478 .513 .520 .533 .482
light .527 .541 .508 .534 .510 .529 .512 .500 .489 .522 .486 .480 .467 .518
inhibited .512 .518 .478 .503 .481 .487 .504 .489 .504 .487 .491 .475 .490
free .488 .482 .522 .497 .519 .513 .496 .511 .496 .513 .509 .525 .510
interesting .520 .508 .491 .496 .492 .505 .503 .493 .506 .488 .492 .501 .497
uninteresting .480 .492 .509 .503 .508 .495 .497 .507 .494 .512 .508 .499 .503
masculine .495 .430 .483 .517 .433 .513 .483 .486 .512 .466 .477 .448 .480
feminine .505 .569 .517 .483 .567 .487 .517 .514 .488 .534 .522 .552 .520
pleasant .498 .491 .483 .505 .469 .497 .490 .478 .497 .493 .465 .496 .490 .499
unpleasant .502 .509 .517 .494 .531 .503 .510 .521 .503 .507 .535 .504 .509 .501
realistic .327 .341 .332 .314 .322 .318 .337 .320 .352 .316 .325 .312 .315 .342
fantastical .673 .659 .668 .686 .678 .682 .663 .680 .647 .685 .675 .688 .685 .658
rugged .474 .467 .477 .480 .453 .488 .489 .477 .484 .478 .469 .452 .463
delicate .525 .532 .523 .519 .547 .512 .511 .523 .515 .522 .531 .547 .537
sharp .509 .557 .483 .506 .582 .517 .459 .474 .478 .490 .474 .506 .503
round .490 .442 .516 .494 .418 .482 .541 .525 .522 .509 .526 .494 .497
simple .499 .520 .509 .511 .524 .503 .486 .497 .507 .507 .489 .511 .519 .517
complex .501 .479 .491 .489 .476 .497 .513 .503 .493 .492 .511 .488 .481 .483
solid .370 .351 .379 .359 .371 .370 .361 .366 .371 .364 .370 .363 .345
nonsolid .629 .648 .621 .641 .629 .630 .639 .633 .628 .636 .629 .636 .655
strong .483 .474 .482 .463 .501 .489 .473 .492 .481 .489 .522 .522 .480
weak .517 .526 .517 .537 .499 .511 .527 .508 .519 .511 .478 .478 .520
structured .361 .364 .346 .359 .360 .355 .359 .355 .376 .350 .350 .347 .360 .361
disorganized .639 .636 .654 .641 .640 .645 .641 .645 .624 .650 .650 .653 .640 .639
tense .501 .516 .494 .499 .515 .497 .496 .503 .494 .499 .499 .511 .497 .491
relaxed .499 .484 .506 .501 .484 .502 .504 .497 .506 .501 .501 .489 .503 .509
Table 13: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - IPA text - Part 1).
Dimension W: a: o: i: e: b d g p t k v D
abrupt .500 .502 .508 .504 .504 .523 .517 .501 .519 .512 .515 .500 .534
continuous .500 .498 .492 .496 .496 .477 .482 .499 .481 .488 .485 .499 .466
active .423 .465 .509 .492 .597 .497 .504 .491 .495 .498 .498 .482
passive .577 .535 .491 .508 .403 .503 .496 .508 .505 .502 .502 .517
beautiful .564 .432 .435 .404 .438 .419 .416 .483 .459 .444 .388
ugly .435 .568 .565 .596 .561 .581 .584 .517 .540 .555 .612
big .636 .596 .613 .533 .547 .567 .570 .507 .501 .500 .635
small .364 .404 .387 .467 .452 .432 .430 .493 .499 .500 .365
dangerous .539 .497 .495 .518 .532 .522 .529 .518 .524 .519 .531
safe .461 .503 .505 .482 .468 .478 .471 .482 .475 .481 .468
exciting .453 .464 .478 .467 .466 .479 .488 .484 .472 .437 .482
calming .547 .536 .522 .533 .534 .520 .512 .516 .527 .562 .518
fast .558 .574 .478 .519 .504 .517 .526 .514 .503 .542 .475
slow .441 .426 .522 .480 .495 .483 .474 .485 .497 .458 .525
good .486 .442 .439 .523 .427 .445 .451 .451 .493 .483 .477 .456
bad .514 .558 .561 .476 .573 .555 .549 .549 .507 .517 .523 .544
happy .458 .382 .422 .401 .459 .464 .474 .540 .511 .501 .424
sad .542 .618 .578 .599 .541 .536 .525 .460 .488 .499 .576
hard .473 .461 .453 .470 .489 .482 .492 .497 .491 .493 .464 .475
soft .527 .538 .547 .530 .511 .517 .508 .503 .509 .507 .536 .525
harsh .276 .292 .332 .258 .287 .300 .344 .274 .305 .305 .291 .317
mellow .724 .708 .668 .742 .713 .700 .656 .726 .695 .695 .709 .683
heavy .490 .546 .509 .514 .507 .539 .532 .530 .496 .499 .499 .627
light .510 .454 .491 .486 .493 .461 .468 .470 .504 .501 .500 .373
inhibited .458 .434 .454 .552 .483 .476 .509 .494 .500 .505 .459
free .542 .566 .546 .448 .517 .524 .490 .506 .500 .495 .541
interesting .457 .506 .496 .514 .516 .511 .503
uninteresting .543 .494 .504 .486 .484 .489 .497
masculine .408 .413 .469 .455 .468 .492 .536 .512 .520
feminine .591 .587 .531 .545 .532 .508 .464 .488 .480
pleasant .497 .446 .438 .455 .430 .480 .477 .482 .488 .480 .471 .450
unpleasant .503 .554 .562 .545 .570 .520 .523 .517 .512 .520 .529 .550
realistic .322 .322 .304 .285 .308 .357 .328 .317 .308 .335 .326 .335 .259
fantastical .678 .678 .696 .715 .692 .643 .672 .683 .692 .665 .674 .666 .741
rugged .468 .459 .498 .471 .479 .488 .471 .488 .494 .489 .490 .483
delicate .532 .541 .502 .529 .521 .512 .529 .512 .505 .511 .510 .517
sharp .460 .486 .523 .504 .499 .499 .513 .512 .511 .519
round .540 .514 .477 .496 .501 .501 .487 .488 .488 .480
simple .511 .522 .528 .545 .503 .523 .520 .512 .499 .510 .489 .484 .515 .519
complex .489 .477 .472 .455 .497 .476 .479 .487 .501 .490 .510 .516 .485 .481
solid .346 .376 .370 .361 .382 .385 .358 .386 .380 .374 .371 .390
nonsolid .654 .624 .630 .639 .618 .615 .642 .614 .620 .626 .629 .610
strong .461 .538 .522 .510 .552 .617 .525 .533 .523 .500 .497 .495 .565
weak .539 .462 .478 .490 .448 .383 .475 .467 .477 .500 .503 .504 .435
structured .417 .386 .354 .372 .337 .326 .347 .353 .363 .358 .335
disorganized .583 .614 .646 .628 .663 .674 .653 .647 .637 .642 .665
tense .485 .552 .529 .543 .524 .502 .521 .515 .518 .506 .513
relaxed .515 .448 .471 .457 .476 .498 .479 .485 .481 .494 .487
Table 14: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - IPA text - Part 2).
Dimension z Z f s S T ç h F m n N N
abrupt .515 .488 .528 .508 .516 .510 .505 .516 .515 .511 .511 .504 .500 .474
continuous .485 .512 .472 .491 .484 .490 .494 .484 .485 .489 .489 .496 .500 .526
active .503 .495 .510 .494 .505 .513 .496 .507 .495 .497 .488 .483 .493 .468
passive .497 .505 .490 .506 .495 .487 .504 .492 .504 .502 .511 .517 .507 .532
beautiful .451 .396 .435 .490 .451 .409 .491 .474 .418 .455 .453 .485 .487 .501
ugly .548 .604 .565 .510 .549 .591 .509 .526 .582 .545 .547 .515 .513 .499
big .507 .559 .499 .473 .561 .617 .459 .510 .491 .497 .486 .507 .535
small .493 .441 .501 .527 .439 .383 .541 .490 .509 .502 .514 .493 .465
dangerous .532 .542 .511 .528 .529 .522 .557 .521 .516 .526 .517 .550 .515 .514
safe .467 .458 .489 .472 .471 .477 .442 .479 .484 .474 .482 .450 .485 .486
exciting .494 .526 .501 .486 .507 .482 .487 .486 .474 .473 .491 .466 .446
calming .505 .474 .498 .513 .493 .517 .512 .514 .526 .527 .509 .533 .554
fast .504 .485 .556 .483 .563 .520 .523 .514 .513 .452 .482 .476 .491
slow .496 .515 .444 .516 .437 .480 .477 .486 .487 .548 .518 .523 .509
good .485 .410 .472 .493 .476 .484 .489 .465 .476 .481 .481 .495 .468
bad .515 .590 .528 .507 .524 .516 .510 .535 .523 .519 .519 .505 .532
happy .437 .508 .516 .472 .484 .410 .529 .563 .492 .488 .549 .514 .436
sad .563 .492 .484 .528 .516 .590 .470 .437 .508 .511 .451 .486 .563
hard .471 .475 .484 .482 .480 .483 .483 .504 .492 .486 .493 .484 .483 .476
soft .528 .525 .515 .518 .520 .517 .517 .496 .508 .513 .506 .516 .517 .524
harsh .321 .264 .305 .291 .295 .283 .304 .321 .262 .269 .286 .263 .276 .268
mellow .679 .736 .695 .709 .705 .717 .696 .679 .738 .731 .714 .737 .724 .732
heavy .523 .524 .497 .478 .514 .545 .483 .469 .496 .500 .479 .460 .509 .512
light .477 .476 .503 .522 .486 .455 .517 .531 .504 .500 .521 .540 .491 .488
inhibited .478 .500 .481 .504 .448 .577 .498 .507 .455 .523 .504 .494 .497 .483
free .522 .500 .519 .496 .552 .423 .502 .493 .545 .477 .496 .506 .503 .517
interesting .510 .506 .507 .507 .512 .504 .498 .475 .513 .487
uninteresting .490 .494 .492 .493 .488 .496 .501 .525 .487 .513
masculine .414 .459 .475 .533 .491 .491 .484 .478 .530 .523 .484
feminine .586 .540 .525 .467 .509 .509 .515 .522 .470 .477 .516
pleasant .486 .430 .471 .493 .475 .459 .490 .491 .492 .478 .476 .480 .488 .490
unpleasant .513 .570 .529 .507 .525 .541 .510 .509 .507 .522 .524 .520 .512 .510
realistic .297 .309 .337 .321 .320 .322 .297 .331 .284 .336 .337 .321 .327 .361
fantastical .703 .691 .663 .679 .680 .678 .703 .669 .716 .664 .663 .679 .673 .639
rugged .493 .452 .478 .496 .461 .445 .478 .505 .491 .486 .494 .496 .503 .500
delicate .507 .548 .521 .503 .538 .554 .522 .495 .509 .514 .506 .504 .497 .500
sharp .531 .467 .540 .511 .564 .536 .518 .517 .515 .486 .492 .492 .481 .477
round .469 .533 .460 .489 .436 .463 .482 .483 .485 .514 .508 .508 .519 .523
simple .504 .490 .515 .503 .513 .518 .501 .513 .512 .494 .492 .486 .490 .490
complex .495 .510 .485 .496 .487 .482 .499 .487 .488 .506 .508 .514 .510 .509
solid .373 .365 .368 .350 .363 .380 .370 .372 .387 .360 .366 .355 .350 .349
nonsolid .627 .635 .632 .650 .637 .619 .630 .627 .613 .640 .634 .644 .650 .651
strong .534 .523 .506 .479 .538 .525 .488 .494 .478 .477 .487 .472 .496 .497
weak .466 .477 .494 .521 .462 .475 .512 .506 .522 .523 .513 .527 .504 .502
structured .327 .366 .323 .340 .364 .360 .351 .316 .306 .338 .333 .361 .354 .360
disorganized .673 .634 .677 .660 .635 .639 .649 .684 .694 .662 .667 .639 .646 .640
tense .523 .551 .536 .501 .540 .542 .515 .499 .512 .504 .498 .514 .499 .487
relaxed .476 .449 .464 .498 .460 .458 .485 .501 .488 .496 .502 .485 .500 .513
Table 15: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - IPA text - Part 3).
Dimension l R R r j C t͡C P w 4
abrupt .507 .509 .506 .502 .518 .493 .514 .515 .511 .500 .534 .507 .511
continuous .492 .490 .494 .497 .482 .507 .485 .485 .489 .500 .466 .492 .489
active .491 .478 .498 .495 .512 .511 .508 .492 .503 .501 .475 .498 .471
passive .508 .521 .501 .505 .488 .489 .492 .508 .497 .499 .525 .501 .529
beautiful .463 .432 .496 .448 .460 .514 .447 .368 .441 .425 .455 .443 .411
ugly .537 .568 .504 .551 .540 .486 .552 .632 .559 .574 .545 .557 .589
big .484 .554 .508 .506 .468 .587 .505 .527 .503 .480 .520 .537 .491
small .516 .445 .492 .494 .532 .413 .495 .473 .497 .520 .480 .463 .508
dangerous .513 .527 .522 .548 .504 .521 .519 .518 .509 .506 .514 .509
safe .486 .473 .477 .451 .496 .478 .481 .482 .491 .494 .486 .491
exciting .465 .474 .482 .481 .493 .445 .486 .488 .481 .477 .466 .470
calming .535 .526 .517 .518 .507 .555 .514 .512 .519 .523 .533 .530
fast .482 .504 .468 .519 .485 .452 .506 .474 .486 .465 .534 .528 .538
slow .518 .496 .531 .481 .515 .548 .494 .526 .514 .534 .465 .472 .462
good .474 .472 .468 .462 .478 .473 .489 .482 .478 .439 .468 .476 .490
bad .526 .528 .532 .538 .522 .527 .511 .518 .522 .561 .532 .524 .510
happy .500 .528 .492 .491 .482 .476 .476 .446 .453 .385 .482 .506
sad .499 .472 .508 .509 .518 .524 .523 .554 .547 .614 .518 .494
hard .482 .489 .494 .485 .485 .466 .497 .487 .505 .469 .487 .484 .487
soft .518 .511 .505 .515 .515 .534 .503 .513 .495 .531 .513 .516 .513
harsh .287 .275 .281 .278 .296 .277 .297 .323 .284 .296 .275 .298 .285
mellow .713 .725 .719 .722 .704 .723 .703 .677 .716 .704 .725 .702 .715
heavy .475 .528 .487 .496 .470 .470 .500 .517 .503 .510 .492 .516 .488
light .525 .472 .512 .504 .530 .530 .499 .483 .497 .490 .508 .484 .512
inhibited .492 .469 .471 .481 .484 .472 .510 .501 .505 .525 .485 .476 .462
free .508 .530 .529 .519 .515 .527 .490 .499 .495 .475 .515 .524 .538
interesting .507 .484 .513 .504 .503 .528 .512 .493 .509
uninteresting .493 .516 .486 .496 .497 .471 .488 .507 .491
masculine .501 .487 .447 .498 .489 .466 .503 .481 .477 .499
feminine .499 .513 .553 .502 .511 .534 .497 .519 .523 .500
pleasant .482 .494 .499 .482 .484 .506 .476 .465 .465 .458 .483 .484 .513
unpleasant .517 .505 .500 .518 .516 .494 .523 .535 .535 .542 .517 .516 .486
realistic .336 .321 .331 .329 .310 .323 .288 .262 .295 .332 .339 .330 .355
fantastical .665 .679 .668 .671 .690 .677 .712 .738 .705 .668 .661 .670 .645
rugged .487 .494 .509 .478 .487 .489 .487 .508 .486 .506 .503 .486 .481
delicate .513 .506 .490 .522 .512 .510 .513 .492 .514 .493 .497 .514 .518
sharp .481 .500 .478 .501 .525 .466 .533 .495 .528 .491 .537 .525 .550
round .519 .500 .522 .499 .475 .534 .466 .505 .472 .509 .463 .474 .449
simple .495 .499 .495 .508 .500 .510 .462 .496 .471 .439 .502 .511 .536
complex .505 .501 .505 .492 .499 .490 .537 .504 .529 .560 .498 .488 .464
solid .367 .367 .369 .354 .353 .372 .367 .387 .374 .365 .385 .370 .350
nonsolid .633 .632 .631 .646 .647 .628 .633 .613 .626 .635 .615 .630 .650
strong .487 .525 .492 .489 .497 .472 .512 .516 .504 .480 .502 .504 .469
weak .513 .475 .508 .511 .503 .528 .488 .484 .496 .520 .498 .496 .531
structured .338 .366 .364 .350 .346 .356 .383 .330 .382 .350 .342 .338 .360
disorganized .662 .634 .636 .650 .654 .644 .617 .670 .618 .649 .658 .662 .640
tense .497 .509 .496 .507 .511 .488 .545 .509 .542 .526 .519 .513 .517
relaxed .503 .491 .504 .493 .489 .512 .454 .491 .458 .474 .481 .487 .482
Table 16: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - IPA text - Part 4).
Dimension i I e E a 1 W u U o 2 O A
abrupt .517 .525 .526 .522 .537 .518 .513 .513 .557 .516 .534 .524 .520 .532
continuous .483 .475 .474 .478 .463 .482 .486 .487 .443 .484 .466 .476 .480 .468
active .500 .506 .496 .496 .502 .494 .504 .496 .485 .497 .509 .490 .511 .511
passive .500 .494 .504 .504 .498 .506 .496 .504 .515 .503 .491 .510 .489 .489
beautiful .467 .467 .460 .488 .479 .469 .479 .439 .531 .470 .479 .437 .475 .466
ugly .533 .533 .540 .512 .521 .531 .521 .561 .469 .530 .521 .563 .525 .534
big .496 .511 .513 .488 .505 .494 .498 .496 .525 .484 .516 .508 .517 .509
small .504 .489 .487 .512 .495 .506 .502 .504 .475 .516 .484 .492 .483 .491
dangerous .520 .526 .533 .526 .524 .532 .510 .519 .518 .517 .512 .526 .516
safe .480 .474 .467 .474 .476 .468 .490 .481 .482 .483 .488 .474 .484
exciting .497 .515 .499 .507 .487 .491 .490 .494 .489 .503 .493 .505 .499
calming .502 .485 .501 .492 .513 .509 .510 .506 .511 .497 .507 .495 .501
fast .480 .495 .492 .486 .486 .476 .482 .468 .463 .492 .477 .494 .488
slow .520 .505 .508 .514 .514 .524 .518 .532 .537 .508 .523 .506 .512
good .498 .500 .477 .487 .486 .501 .499 .499 .562 .499 .505 .491 .496 .491
bad .502 .500 .523 .513 .514 .499 .501 .501 .438 .501 .495 .509 .504 .509
happy .487 .467 .491 .487 .488 .488 .475 .486 .484 .487 .480 .484 .482
sad .513 .533 .509 .512 .512 .512 .525 .514 .516 .513 .520 .516 .518
hard .495 .499 .503 .505 .501 .497 .487 .493 .494 .487 .499 .495 .500
soft .504 .501 .497 .494 .499 .503 .513 .507 .506 .513 .500 .505 .500
harsh .299 .315 .309 .320 .294 .315 .286 .314 .319 .301 .308 .329 .306 .311
mellow .701 .685 .691 .680 .706 .685 .714 .686 .681 .699 .692 .671 .694 .689
heavy .511 .504 .519 .506 .520 .510 .516 .521 .528 .517 .515 .526 .530 .521
light .489 .496 .481 .494 .480 .490 .484 .479 .472 .483 .485 .474 .470 .479
inhibited .504 .491 .502 .492 .510 .500 .507 .493 .501 .496 .499 .489 .492
free .496 .509 .498 .508 .490 .500 .493 .507 .499 .504 .501 .511 .508
interesting .504 .500 .519 .510 .507 .498 .496 .510 .507 .503 .511 .504
uninteresting .496 .500 .481 .490 .493 .502 .504 .490 .493 .497 .489 .496
masculine .501 .486 .494 .529 .496 .517 .496 .506 .503 .493 .513 .494 .497
feminine .498 .514 .506 .471 .504 .483 .504 .494 .497 .507 .487 .506 .503
pleasant .492 .502 .481 .494 .493 .487 .492 .482 .503 .489 .504 .482 .494 .497
unpleasant .508 .498 .519 .506 .507 .513 .508 .518 .497 .511 .496 .518 .506 .503
realistic .353 .339 .352 .357 .333 .341 .336 .342 .365 .342 .353 .342 .333 .335
fantastical .647 .661 .648 .643 .667 .659 .664 .658 .635 .658 .647 .658 .667 .665
rugged .465 .485 .470 .473 .474 .475 .474 .479 .478 .488 .479 .485 .485
delicate .535 .515 .530 .527 .526 .525 .526 .521 .522 .512 .521 .515 .515
sharp .522 .533 .515 .528 .522 .513 .506 .498 .499 .523 .523 .519 .526
round .478 .467 .485 .472 .478 .487 .494 .502 .501 .476 .477 .481 .474
simple .513 .514 .508 .506 .510 .507 .509 .506 .525 .507 .510 .494 .508 .513
complex .487 .486 .492 .494 .490 .493 .491 .494 .475 .493 .490 .506 .492 .487
solid .376 .362 .376 .384 .375 .365 .375 .378 .382 .362 .379 .380 .370
nonsolid .624 .638 .624 .616 .625 .635 .625 .622 .618 .638 .621 .620 .630
strong .500 .496 .497 .504 .501 .500 .498 .502 .503 .502 .505 .510 .511
weak .500 .504 .503 .496 .499 .500 .502 .498 .497 .498 .495 .490 .489
structured .343 .350 .343 .355 .339 .346 .348 .354 .369 .346 .336 .342 .345 .343
disorganized .657 .650 .657 .645 .661 .654 .652 .646 .631 .654 .664 .658 .655 .657
tense .512 .512 .516 .517 .503 .517 .502 .512 .520 .511 .492 .520 .507 .504
relaxed .488 .488 .484 .483 .497 .483 .498 .487 .480 .489 .508 .480 .493 .496
Table 17: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - Audio - Part 1).
Dimension 5 @ W: a: o: i: b d g p t c k v
abrupt .530 .536 .515 .541 .546 .555 .516 .518 .515 .518 .524 .514 .521 .505
continuous .470 .464 .485 .459 .454 .445 .484 .482 .485 .482 .476 .486 .479 .495
active .494 .503 .508 .516 .508 .504 .507 .512 .506 .504 .497 .508 .502 .509
passive .506 .497 .492 .484 .492 .496 .493 .488 .494 .496 .503 .492 .498 .491
beautiful .464 .471 .488 .444 .469 .455 .479 .491 .469 .505 .463 .430
ugly .536 .529 .512 .556 .531 .545 .521 .509 .531 .495 .537 .570
big .486 .505 .468 .513 .512 .520 .505 .504 .501 .494 .498 .501
small .514 .495 .532 .487 .488 .480 .495 .496 .499 .506 .501 .499
dangerous .512 .526 .487 .532 .543 .468 .517 .521 .516 .497 .522 .512 .518 .503
safe .488 .473 .513 .468 .457 .532 .483 .479 .484 .503 .478 .488 .482 .497
exciting .487 .488 .514 .484 .430 .508 .489 .485 .493 .496 .490 .505 .498 .496
calming .512 .512 .486 .516 .570 .492 .510 .515 .507 .504 .510 .495 .502 .504
fast .479 .477 .506 .482 .490 .483 .484 .490 .482 .480 .487 .514
slow .521 .523 .494 .518 .510 .517 .516 .510 .517 .520 .513 .486
good .480 .493 .514 .499 .479 .454 .487 .489 .504 .499 .495 .507 .496 .455
bad .520 .507 .486 .501 .521 .546 .513 .511 .496 .501 .505 .493 .504 .545
happy .479 .481 .496 .492 .480 .522 .488 .499 .488 .475 .481
sad .521 .518 .504 .508 .520 .478 .512 .501 .512 .525 .519
hard .507 .500 .478 .504 .498 .503 .488 .492 .493 .494 .499 .482 .492 .491
soft .493 .500 .522 .496 .502 .497 .512 .508 .507 .506 .501 .518 .508 .509
harsh .324 .317 .298 .298 .275 .298 .285 .286 .292 .298 .315 .297 .307 .276
mellow .676 .683 .703 .702 .725 .702 .715 .714 .708 .702 .685 .703 .693 .724
heavy .509 .521 .515 .531 .535 .493 .523 .520 .515 .517 .521 .510 .515 .528
light .491 .479 .485 .469 .465 .507 .477 .480 .485 .483 .479 .490 .485 .472
inhibited .501 .500 .504 .503 .505 .485 .505 .507 .499 .494 .502 .502 .502 .495
free .499 .500 .496 .497 .495 .515 .495 .492 .501 .506 .498 .497 .498 .505
interesting .507 .503 .522 .503 .491 .496 .504 .504 .525 .504
uninteresting .493 .497 .478 .497 .509 .504 .496 .496 .475 .496
masculine .519 .461 .514 .508 .501 .500 .504 .505 .507 .482 .516
feminine .481 .539 .486 .492 .499 .500 .496 .495 .493 .518 .484
pleasant .488 .496 .506 .461 .493 .483 .497 .496 .498 .498 .492 .513 .491 .497
unpleasant .512 .504 .494 .539 .507 .517 .503 .504 .502 .502 .508 .487 .508 .503
realistic .346 .332 .350 .343 .333 .350 .352 .356 .358 .358 .347 .342 .353 .348
fantastical .655 .668 .650 .657 .667 .650 .648 .644 .642 .642 .653 .658 .648 .652
rugged .478 .489 .466 .475 .482 .457 .471 .459 .477 .479 .476 .474 .476 .473
delicate .522 .511 .534 .525 .518 .543 .529 .541 .523 .521 .524 .526 .524 .527
sharp .528 .512 .517 .520 .539 .504 .507 .510 .511 .518 .535 .519 .521
round .471 .487 .483 .480 .461 .496 .493 .490 .489 .482 .465 .481 .479
simple .505 .511 .521 .521 .516 .517 .520 .518 .517 .519 .511 .514 .511 .517
complex .495 .489 .479 .479 .484 .483 .480 .482 .483 .481 .489 .486 .489 .483
solid .382 .374 .371 .375 .390 .364 .371 .371 .370 .372 .378 .374 .377 .416
nonsolid .618 .626 .629 .625 .610 .636 .629 .629 .630 .628 .621 .626 .623 .584
strong .506 .507 .510 .512 .493 .514 .504 .502 .506 .502 .506 .501 .500 .508
weak .494 .493 .490 .488 .507 .486 .496 .498 .494 .498 .494 .499 .500 .492
structured .349 .341 .347 .351 .318 .354 .339 .344 .344 .347 .349 .351 .344 .335
disorganized .651 .659 .653 .649 .682 .646 .661 .656 .656 .653 .651 .649 .656 .665
tense .511 .512 .503 .509 .515 .506 .496 .503 .495 .502 .513 .511 .506 .492
relaxed .489 .488 .497 .491 .485 .494 .504 .497 .505 .498 .487 .489 .494 .508
Table 18: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - Audio - Part 2).
Dimension z K f s S T h F m n N N l
abrupt .514 .534 .514 .505 .525 .521 .514 .513 .511 .512 .500 .519 .522 .518
continuous .485 .466 .486 .495 .475 .479 .486 .487 .489 .488 .500 .481 .477 .482
active .508 .502 .506 .499 .508 .516 .505 .503 .505 .505 .502 .500 .520 .505
passive .492 .498 .494 .501 .492 .484 .495 .497 .495 .495 .498 .500 .480 .495
beautiful .464 .442 .498 .487 .495 .474 .488 .529 .472 .473 .491 .482 .505 .484
ugly .535 .558 .502 .513 .504 .526 .512 .471 .528 .527 .509 .518 .495 .516
big .484 .506 .509 .484 .512 .478 .504 .498 .492 .501 .495 .500 .506
small .516 .494 .490 .516 .488 .521 .495 .502 .508 .499 .505 .500 .494
dangerous .517 .536 .506 .499 .508 .502 .506 .482 .515 .525 .501 .518 .521 .509
safe .483 .464 .494 .501 .492 .498 .494 .517 .485 .475 .499 .482 .478 .491
exciting .500 .495 .497 .498 .495 .499 .498 .494 .493 .497 .496 .501 .496 .496
calming .500 .505 .503 .502 .505 .501 .502 .506 .507 .503 .504 .499 .504 .504
fast .484 .495 .501 .477 .508 .482 .489 .462 .477 .487 .471 .485 .496 .479
slow .516 .505 .499 .523 .492 .518 .511 .538 .523 .513 .529 .515 .504 .521
good .487 .474 .499 .503 .493 .505 .491 .500 .504 .518 .502 .508 .494
bad .512 .526 .501 .497 .507 .495 .509 .500 .496 .482 .498 .492 .506
happy .484 .490 .472 .484 .478 .478 .502 .499 .490 .484 .496 .489 .493 .489
sad .516 .510 .528 .515 .522 .522 .497 .501 .510 .516 .504 .511 .507 .511
hard .489 .499 .485 .486 .486 .491 .493 .472 .494 .497 .485 .497 .489 .493
soft .511 .501 .515 .514 .514 .509 .507 .527 .506 .503 .515 .503 .510 .507
harsh .298 .290 .298 .305 .301 .280 .298 .291 .294 .288 .277 .301 .289 .303
mellow .702 .710 .702 .695 .699 .720 .702 .709 .706 .712 .723 .699 .711 .697
heavy .503 .532 .514 .508 .517 .510 .506 .498 .524 .520 .496 .518 .515 .517
light .497 .468 .486 .492 .483 .490 .494 .502 .476 .480 .504 .482 .485 .483
inhibited .504 .474 .493 .498 .492 .511 .498 .499 .500 .505 .506 .506 .501 .499
free .495 .526 .507 .502 .508 .489 .502 .500 .500 .495 .494 .494 .499 .501
interesting .503 .514 .514 .504 .519 .506 .501 .498 .501 .495 .518 .500
uninteresting .497 .486 .486 .496 .481 .494 .499 .502 .499 .505 .482 .500
masculine .466 .498 .512 .515 .490 .475 .504 .500 .506 .518 .501 .509
feminine .534 .502 .488 .485 .510 .525 .496 .500 .494 .481 .498 .491
pleasant .489 .495 .499 .497 .494 .510 .495 .490 .488 .492 .503 .492 .500 .501
unpleasant .511 .505 .501 .503 .506 .489 .505 .510 .512 .508 .497 .508 .500 .499
realistic .350 .337 .359 .357 .357 .375 .365 .342 .352 .353 .344 .353 .359 .345
fantastical .650 .663 .641 .643 .643 .625 .635 .658 .648 .647 .656 .648 .641 .655
rugged .480 .484 .480 .471 .457 .460 .472 .465 .472 .474 .459 .478 .470 .477
delicate .520 .516 .520 .529 .543 .540 .528 .535 .528 .526 .541 .522 .530 .523
sharp .528 .518 .520 .523 .534 .514 .515 .499 .500 .518 .500 .512 .512 .513
round .472 .482 .480 .477 .466 .485 .485 .501 .500 .482 .500 .488 .488 .486
simple .517 .507 .520 .515 .515 .532 .517 .520 .511 .509 .519 .512 .524 .512
complex .483 .493 .480 .485 .485 .468 .483 .480 .489 .491 .481 .488 .476 .488
solid .358 .373 .364 .362 .362 .383 .369 .357 .363 .360 .354 .370 .374 .367
nonsolid .642 .627 .636 .638 .638 .617 .631 .643 .637 .640 .646 .630 .626 .633
strong .504 .506 .502 .501 .515 .503 .497 .487 .502 .504 .499 .507 .504 .501
weak .496 .494 .498 .499 .485 .497 .503 .513 .498 .496 .501 .493 .496 .499
structured .328 .340 .337 .343 .331 .356 .338 .343 .337 .334 .346 .341 .349 .341
disorganized .672 .660 .663 .657 .669 .644 .662 .657 .663 .666 .654 .659 .651 .659
tense .507 .508 .493 .506 .502 .502 .495 .497 .504 .510 .504 .515 .499 .497
relaxed .493 .492 .507 .494 .498 .498 .505 .503 .496 .490 .496 .485 .501 .503
Table 19: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - Audio - Part 3).
Dimension l (retroflex) r R j C P w 4
abrupt .523 .525 .513 .516 .513 .514 .518 .523 .528
continuous .477 .475 .487 .484 .487 .486 .482 .477 .472
active .498 .508 .497 .504 .511 .510 .514 .504 .491
passive .502 .492 .503 .496 .489 .490 .486 .496 .509
beautiful .462 .483 .488 .490 .468 .449 .499 .466 .431
ugly .538 .517 .512 .510 .532 .551 .501 .534 .569
big .496 .519 .479 .495 .498 .492 .505 .500 .501
small .504 .481 .521 .505 .502 .507 .495 .500 .499
dangerous .527 .530 .541 .515 .506 .520 .512 .505 .515
safe .472 .470 .459 .485 .494 .480 .488 .495 .485
exciting .501 .496 .482 .487 .495 .484 .500 .495
calming .499 .504 .518 .513 .505 .515 .500 .505
fast .474 .485 .459 .473 .489 .450 .506 .484 .451
slow .526 .515 .541 .527 .511 .550 .494 .516 .549
good .493 .497 .487 .501 .488 .482 .496 .494 .414
bad .507 .503 .513 .498 .512 .518 .504 .506 .586
happy .484 .483 .475 .487 .497 .515 .491 .500
sad .516 .517 .525 .513 .503 .485 .508 .500
hard .502 .499 .490 .484 .493 .483 .485 .488 .494
soft .497 .501 .510 .515 .507 .517 .515 .512 .506
harsh .308 .316 .306 .288 .309 .303 .278 .303 .258
mellow .692 .684 .694 .712 .691 .697 .722 .697 .742
heavy .515 .517 .500 .508 .529 .508 .507 .505
light .485 .483 .499 .492 .471 .492 .493 .495
inhibited .499 .479 .496 .494 .505 .509 .495 .496 .479
free .500 .521 .504 .506 .495 .490 .505 .504 .521
interesting .501 .513 .463 .504 .504 .507
uninteresting .499 .487 .537 .496 .496 .493
masculine .515 .508 .476 .486 .492 .498 .503
feminine .485 .492 .523 .514 .508 .502 .497
pleasant .488 .499 .498 .500 .490 .464 .507 .494 .506
unpleasant .512 .501 .502 .500 .510 .536 .493 .506 .494
realistic .347 .349 .321 .337 .346 .333 .369 .354 .342
fantastical .653 .651 .679 .663 .654 .667 .631 .646 .658
rugged .473 .494 .470 .466 .472 .492 .468 .472 .458
delicate .527 .506 .530 .534 .528 .508 .532 .527 .542
sharp .515 .523 .501 .513 .546 .520 .525 .513 .521
round .485 .477 .499 .487 .454 .480 .475 .487 .479
simple .501 .510 .509 .520 .516 .499 .521 .515 .515
complex .499 .490 .491 .480 .484 .501 .479 .485 .485
solid .378 .374 .358 .364 .376 .371 .373 .374 .395
nonsolid .622 .626 .642 .636 .624 .629 .627 .626 .604
strong .501 .509 .497 .500 .513 .477 .506 .497 .489
weak .499 .491 .503 .500 .487 .523 .494 .503 .511
structured .343 .345 .341 .337 .350 .335 .345 .338 .343
disorganized .657 .655 .659 .663 .650 .665 .655 .662 .657
tense .509 .502 .505 .492 .529 .512 .500 .494 .511
relaxed .490 .498 .495 .508 .471 .488 .500 .506 .489
Table 20: Phoneme-Semantic Feature Attention Fraction Scores (Natural words - Audio - Part 4).
Dimension i ej A ow b d g p t k v D z f
abrupt .513 .527 .514 .513 .527 .533 .527 .506 .497 .511 .465 .460 .491 .521
continuous .487 .473 .486 .486 .473 .467 .472 .494 .503 .489 .535 .540 .509 .479
active .498 .498 .485 .487 .498 .509 .497 .495 .500 .506 .504 .496 .502 .498
passive .501 .501 .514 .512 .501 .491 .502 .504 .500 .494 .496 .504 .497 .502
beautiful .522 .534 .529 .511 .434 .410 .408 .543
ugly .478 .465 .471 .489 .565 .590 .592 .456
big .463 .439 .586 .567 .550 .555 .547 .521 .508 .505 .556 .527 .541 .491
small .537 .561 .414 .433 .450 .445 .453 .479 .492 .494 .444 .473 .459 .509
dangerous .513 .513 .510 .510 .522 .528 .531 .527 .530 .526
safe .486 .487 .490 .489 .478 .471 .469 .472 .470 .474
exciting .509 .524 .465 .480 .498 .499 .496 .521 .529 .531 .533 .482 .527 .512
calming .491 .475 .534 .519 .501 .501 .503 .478 .471 .469 .467 .518 .473 .488
fast .548 .551 .453 .438 .493 .484 .490 .555 .555 .547 .556 .537 .540 .494
slow .451 .449 .547 .562 .507 .516 .510 .445 .445 .453 .444 .463 .460 .506
happy .567 .577 .412 .425 .485 .476 .480 .495 .485 .487 .484 .500 .476 .492
sad .432 .423 .588 .574 .514 .524 .520 .505 .514 .513 .516 .500 .524 .508
hard .482 .484 .488 .467 .494 .488 .485 .534 .518 .519 .492 .525 .485
soft .517 .516 .512 .532 .506 .512 .515 .466 .482 .480 .507 .475 .515
harsh .262 .281 .267 .288 .283 .284 .258 .350 .348 .337 .350 .342 .340 .288
mellow .738 .719 .733 .712 .717 .716 .742 .650 .652 .663 .650 .658 .660 .712
heavy .452 .443 .543 .539 .560 .553 .541 .511 .506 .512 .556 .527 .537 .480
light .548 .557 .457 .461 .440 .447 .459 .489 .494 .488 .444 .473 .463 .520
inhibited .497 .474 .468 .467 .560 .564 .561
free .503 .525 .532 .533 .440 .436 .439
masculine .472 .472 .497 .477 .506 .487 .505 .504 .500 .493 .497 .481 .500 .478
feminine .528 .528 .503 .523 .494 .513 .495 .496 .500 .506 .503 .519 .499 .521
realistic .335 .328 .366 .366 .379 .400 .348 .340 .325 .290 .381 .378 .384 .362
fantastical .665 .672 .634 .634 .621 .600 .652 .660 .675 .710 .619 .622 .616 .638
rugged .464 .462 .478 .478 .505 .487 .499 .460 .439 .460 .416 .401 .420 .485
delicate .536 .538 .522 .521 .494 .513 .500 .540 .561 .540 .584 .598 .580 .514
sharp .537 .541 .486 .482 .498 .492 .493 .542 .543 .543 .554 .508 .554 .523
round .462 .459 .513 .518 .501 .507 .506 .458 .456 .456 .445 .492 .445 .477
simple .519 .520 .512 .526 .522 .519 .499 .519 .514 .504 .433 .434 .514
complex .481 .480 .488 .473 .478 .481 .500 .481 .485 .495 .567 .565 .486
solid .358 .360 .362 .364 .365 .357 .371 .422 .428 .466 .362 .359 .356 .354
nonsolid .642 .640 .638 .636 .635 .643 .629 .578 .571 .534 .638 .641 .643 .646
strong .495 .482 .486 .497 .556 .563 .571
weak .505 .518 .514 .503 .444 .437 .429
tense .505 .513 .500 .501 .566 .578 .556
relaxed .494 .487 .500 .499 .434 .421 .444
Table 21: Phoneme-Semantic Feature Attention Fraction Scores (Constructed words - IPA text - Part 1).
Dimension s S m n l
abrupt .516 .513 .515 .529 .515
continuous .484 .487 .485 .471 .485
active .494 .498 .481 .500 .489
passive .505 .502 .519 .500 .511
beautiful .556 .529 .551 .553 .537
ugly .444 .471 .449 .447 .463
big .478 .481 .463 .465 .462
small .522 .519 .537 .535 .538
dangerous .456 .455 .477
safe .544 .545 .523
exciting .505 .511 .480 .486 .481
calming .495 .489 .519 .514 .519
fast .478 .477 .439 .447 .454
slow .522 .523 .560 .553 .545
happy .479 .483 .488 .495 .483
sad .521 .517 .511 .504 .517
hard .481 .476 .488 .483 .483
soft .518 .524 .512 .517 .517
harsh .283 .279 .267 .275 .280
mellow .717 .721 .733 .725 .720
heavy .465 .460 .444 .438 .441
light .534 .540 .556 .562 .559
inhibited .444 .450 .439
free .556 .550 .561
masculine .488 .483 .468 .498 .500
feminine .512 .517 .532 .501 .500
realistic .379 .351 .366 .384 .377
fantastical .621 .649 .634 .616 .623
rugged .488 .476 .482 .494 .482
delicate .511 .523 .518 .505 .518
sharp .515 .519 .486 .473 .483
round .485 .481 .514 .527 .517
simple .522 .511 .515 .520 .524
complex .477 .489 .485 .480 .475
solid .334 .359 .359 .345 .367
nonsolid .666 .641 .641 .655 .633
strong .454 .470 .476
weak .546 .530 .524
tense .486 .475 .481
relaxed .514 .524 .519
Table 22: Phoneme-Semantic Feature Attention Fraction Scores (Constructed words - IPA text - Part 2).
Dimension i A b d g p t k v z f s S m
abrupt .522 .539 .531 .521 .529 .505 .513 .521 .512 .520 .526 .523 .529 .526
continuous .478 .461 .469 .479 .471 .495 .486 .479 .488 .480 .474 .477 .471 .474
active .507 .510 .516 .511 .509 .515 .509 .510 .513 .506 .506 .513 .508 .515
passive .493 .490 .484 .489 .491 .485 .491 .490 .487 .494 .494 .487 .491 .485
beautiful .511 .505 .477 .474 .477 .522 .523 .518 .517
ugly .489 .495 .523 .526 .523 .478 .477 .482 .483
big .499 .536 .521 .514 .517 .502 .501 .498 .516 .507 .503 .493 .499 .498
small .501 .464 .479 .486 .483 .498 .499 .502 .484 .493 .497 .507 .501 .502
dangerous .509 .517 .502 .507 .504 .511 .514 .552
safe .491 .483 .498 .493 .495 .489 .486 .448
exciting .510 .498 .504 .506 .509 .496 .520 .504 .511 .521 .515 .511 .515 .499
calming .490 .502 .496 .494 .491 .504 .480 .496 .489 .479 .485 .489 .485 .501
fast .525 .465 .492 .490 .491 .519 .536 .530 .519 .528 .495 .488 .481 .467
slow .475 .535 .508 .510 .509 .481 .464 .470 .481 .472 .505 .512 .519 .533
happy .508 .476 .506 .499 .498 .503 .503 .506 .505 .496 .508 .487 .497 .499
sad .492 .524 .494 .500 .502 .497 .497 .494 .495 .504 .491 .513 .503 .501
hard .490 .489 .489 .488 .484 .504 .510 .520 .520 .519 .467 .469 .473 .487
soft .510 .511 .511 .512 .516 .496 .490 .480 .480 .481 .533 .531 .527 .513
harsh .282 .297 .275 .278 .272 .319 .319 .314 .315 .319 .280 .267 .277 .265
mellow .718 .703 .725 .722 .728 .681 .681 .686 .685 .681 .721 .733 .723 .735
heavy .475 .522 .519 .519 .518 .497 .496 .503 .526 .519 .485 .480 .484 .463
light .525 .478 .481 .481 .482 .503 .504 .497 .474 .481 .515 .520 .516 .537
inhibited .475 .472 .473 .493 .494 .461
free .525 .528 .527 .506 .506 .539
masculine .487 .509 .502 .500 .504 .502 .498 .507 .500 .503 .492 .501 .492 .492
feminine .513 .491 .498 .500 .496 .498 .502 .493 .500 .497 .508 .499 .508 .508
realistic .366 .356 .366 .372 .360 .361 .368 .365 .363 .373 .354 .374 .367 .364
fantastical .634 .644 .634 .628 .640 .639 .632 .635 .637 .627 .646 .626 .633 .636
rugged .480 .498 .508 .491 .496 .476 .476 .481 .474 .481 .479 .504 .475 .490
delicate .520 .502 .492 .509 .504 .524 .524 .519 .526 .519 .521 .496 .525 .510
sharp .535 .505 .507 .503 .495 .533 .537 .533 .534 .541 .524 .524 .530 .498
round .465 .495 .493 .497 .504 .467 .463 .467 .466 .458 .476 .476 .470 .501
simple .517 .515 .515 .518 .512 .518 .516 .511 .471 .466 .516 .519 .515 .516
complex .483 .485 .485 .482 .488 .482 .484 .489 .529 .534 .484 .481 .485 .484
solid .356 .367 .350 .352 .351 .387 .392 .392 .359 .349 .359 .349 .345 .351
nonsolid .644 .633 .650 .648 .649 .613 .608 .608 .641 .651 .641 .651 .655 .649
strong .514 .500 .516 .521 .494
weak .486 .500 .484 .479 .506
tense .513 .510 .500 .504 .503 .511
relaxed .487 .490 .499 .496 .497 .489
Table 23: Phoneme-Semantic Feature Attention Fraction Scores (Constructed words - Audio - Part 1).
Dimension n l
abrupt .522 .524
continuous .478 .476
active .513 .518
passive .487 .482
beautiful .530 .514
ugly .470 .486
big .499 .501
small .501 .499
dangerous .558 .535
safe .442 .465
exciting .503 .499
calming .497 .500
fast .456 .462
slow .544 .538
happy .496 .495
sad .504 .505
hard .482 .479
soft .518 .521
harsh .259 .272
mellow .741 .728
heavy .463 .462
light .537 .538
inhibited .444 .456
free .556 .544
masculine .495 .496
feminine .505 .504
realistic .367 .356
fantastical .633 .644
rugged .498 .487
delicate .502 .513
sharp .489 .497
round .511 .503
simple .519 .521
complex .481 .479
solid .349 .354
nonsolid .651 .646
strong .491 .492
weak .509 .508
tense .511 .493
relaxed .489 .507
Table 24: Phoneme-Semantic Feature Attention Fraction Scores (Constructed words - Audio - Part 2).
We also calculate the average attention fraction scores by layer in Qwen2.5-Omni-7B as shown in Table 25.
Layer Natural Constructed
IPA Audio IPA Audio
0 0.5022 0.4974 0.5151 0.4950
1 0.5047 0.5000 0.5048 0.4901
2 0.5060 0.5023 0.5109 0.5000
3 0.5054 0.5002 0.5147 0.4982
4 0.5020 0.5036 0.5100 0.5036
5 0.5090 0.5018 0.5228 0.5052
6 0.5026 0.4985 0.5022 0.4845
7 0.5050 0.5007 0.5394 0.5152
8 0.5087 0.5049 0.5395 0.5185
9 0.5121 0.5076 0.5110 0.5045
10 0.5112 0.5030 0.5253 0.5083
11 0.5096 0.5022 0.5311 0.5088
12 0.5064 0.4957 0.5409 0.5036
13 0.5061 0.4964 0.5270 0.5049
14 0.5065 0.4962 0.5269 0.5111
15 0.5077 0.5046 0.5282 0.5048
16 0.5140 0.5022 0.5249 0.4974
17 0.5075 0.5011 0.5180 0.5023
18 0.5065 0.5058 0.5291 0.5144
19 0.5092 0.5002 0.5062 0.5019
20 0.5055 0.5056 0.5165 0.5104
21 0.5091 0.5016 0.5186 0.5092
22 0.5119 0.5002 0.5285 0.5095
23 0.5082 0.4985 0.5229 0.5148
24 0.5049 0.4991 0.5427 0.5116
25 0.5077 0.4998 0.5337 0.5101
26 0.5104 0.4983 0.5324 0.5076
27 0.5056 0.4987 0.5101 0.5043
Average 0.5073 0.5009 0.5226 0.5053
Table 25: Layer-wise attention fraction scores for IPA and audio input types across natural and constructed word groups in Qwen2.5-Omni-7B.
Appendix G Prompts
Table 26 shows an automatic semantic dimension annotation prompt for LLMs. Table 27 represents detailed prompts for the semantic dimension experiments.
You are a professional linguistic annotator.
Please read a {language} mimetic word and its meaning, and decide which semantic feature best describes the word's meaning.
[WORD]
{word}
[MEANING]
{meaning}
[SEMANTIC DIMENSION]
{feature1} vs. {feature2}
[OPTIONS]
1: {feature1}
2: {feature2}
3: Neither
Answer with the number only. (1-3)
Table 26: Prompt for automatic semantic dimension annotation.
Given a [WORD], which semantic feature best describes the word based on auditory impression?
[WORD]
{word}
[SEMANTIC DIMENSION]
{feature1} vs. {feature2}
[OPTIONS]
1: {feature1}
2: {feature2}
Answer with the number only. (1-2)
(a) Prompt for original text words.
Given an IPA [WORD], which semantic feature best describes the word based on auditory impression?
[WORD]
{word}
[SEMANTIC DIMENSION]
{feature1} vs. {feature2}
[OPTIONS]
1: {feature1}
2: {feature2}
Answer with the number only. (1-2)
(b) Prompt for IPA text words.
Given a spoken [WORD], which semantic feature best describes the word based on auditory impression?
[WORD]
{audio}
[SEMANTIC DIMENSION]
{feature1} vs. {feature2}
[OPTIONS]
1: {feature1}
2: {feature2}
Answer with the number only. (1-2)
(c) Prompt for spoken audio words.
Table 27: Prompts for the semantic dimension prediction experiment by input type.

### v5

@@
 We conduct quantitative evaluations for up to 25 semantic dimensions and examine layer-wise attention fraction scores to identify how phonemes and meanings are related within the models.
-Sound symbolism, which suggests that phonetic sounds and their meanings have a significant correlation, presents a cognitively grounded exception to the principle of linguistic arbitrariness (hinton2006sound; dingemanse2015arbitrariness).
-For instance, when people are presented with two images of sharp and round shapes called ``kiki'' and ``bouba'', the overwhelming majority of participants consistently match ``kiki'' with pointed shapes and ``bouba'' with round shapes, regardless of their cultural background (kohler1967gestalt; ramachandran2001synaesthesia).
-This iconicity effect illustrates the universal sensory associations of humans that intuitively link certain phonetic sounds with certain semantic features, which facilitate early-childhood or second-language acquisition (imai2014sound), and commercially nuanced brand naming (yorkston2004sound).
-Recent advancements in Multimodal Large Language Models (MLLMs) that incorporate audio-modality input into an integrated representation space (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport) shed light on new possibilities to systematically analyze human-like phonetic iconicity of language models.
+Sound symbolism, which suggests that phonetic sounds and their meanings have a significant correlation, presents a cognitively grounded exception to the principle of linguistic arbitrariness (Hinton and others 2006; Dingemanse et al.
+2015).
+For instance, when people are presented with two images of sharp and round shapes called ``kiki'' and ``bouba'', the overwhelming majority of participants consistently match ``kiki'' with pointed shapes and ``bouba'' with round shapes, regardless of their cultural background (Köhler 1967; Ramachandran and Hubbard 2001).
+This iconicity effect illustrates the universal sensory associations of humans that intuitively link certain phonetic sounds with certain semantic features, which facilitate early-childhood or second-language acquisition (Imai and Kita 2014), and commercially nuanced brand naming (Yorkston and Menon 2004).
+Recent advancements in Multimodal Large Language Models (MLLMs) that incorporate audio-modality input into an integrated representation space (OpenAI et al.
+2024; Xu et al.
+2025) shed light on new possibilities to systematically analyze human-like phonetic iconicity of language models.
 With MLLMs as a novel test bed, we formulate the following two key research questions:
@@
 Mimetic words provide a compelling probe to address these subjects.
-Represented by onomatopoeias and ideophones [^1] In this paper, “onomatopoeia” and “ideophone” refer to words that depict sounds and non-auditory sensory imagery, respectively., mimetic words refer to words in which the non-arbitrary association of sound form and meaning is prominent (akita2019ideophones), such as ``boom'' or ``whizz.'' We construct LEX-ICON, a dataset of natural mimetic words and constructed pseudo-words designed to maximize sound-symbolic effect.
+Represented by onomatopoeias and ideophones [^1] In this paper, “onomatopoeia” and “ideophone” refer to words that depict sounds and non-auditory sensory imagery, respectively., mimetic words refer to words in which the non-arbitrary association of sound form and meaning is prominent (Akita and Pardeshi 2019), such as ``boom'' or ``whizz.'' We construct LEX-ICON, a dataset of natural mimetic words and constructed pseudo-words designed to maximize sound-symbolic effect.
 With LEX-ICON, we apply the semantic dimension paradigm, where word meanings are projected onto binary scales (e.g., fast vs.
@@
 2.1 Sound Symbolism in Linguistics
-Early modern linguists question the arbitrary relationship between signifier and signified, suggesting the phenomenon of intuitive phonetic symbolism in humans (usnadze1924experimenteller; sapir1929study; kohler1967gestalt).
-sapir1929study's experiment has exhibited that people who are presented with unfamiliar object names ``mil'' and ``mal'' and are asked to estimate the size of the objects respond that the latter is larger than the former (parise2012audiovisual).
-Recent studies continue to observe these phenomena (hinton2006sound; lockwood2015iconicity; cwiek2022bouba), extending them by scaling to a large amount of data (thompson2021articulatory; winter2024iconicity) or measuring the exact semantic dimensions associated with each phoneme (monaghan2019sound; sidhu2022higher; sidhu2025sound).
-For the mimetic words, experimental research has also demonstrated that people can infer a certain degree of meaning from the words in languages besides their mother tongue (shinohara2010cross; dingemanse2016sound).
+Early modern linguists question the arbitrary relationship between signifier and signified, suggesting the phenomenon of intuitive phonetic symbolism in humans (Usnadze 1924; Sapir 1929; Köhler 1967).
+Sapir (1929)'s experiment has exhibited that people who are presented with unfamiliar object names ``mil'' and ``mal'' and are asked to estimate the size of the objects respond that the latter is larger than the former (Parise and Spence 2012).
+Recent studies continue to observe these phenomena (Hinton and others 2006; Lockwood and Dingemanse 2015; Ćwiek et al.
+2022), extending them by scaling to a large amount of data (Thompson et al.
+2021; Winter et al.
+2024) or measuring the exact semantic dimensions associated with each phoneme (Monaghan and Fletcher 2019; Sidhu et al.
+2022; Sidhu 2025).
+For the mimetic words, experimental research has also demonstrated that people can infer a certain degree of meaning from the words in languages besides their mother tongue (Shinohara and Kawahara 2010; Dingemanse et al.
+2016).
 2.2 Phonetic Iconicity for LLMs
-Iconicity tasks for language models have been used as a means of assessing whether the models have human-like phonetic intuition (cai-etal-2024-large; duan2024hlbbenchmarkingllmshumanlikeness).
-Early studies focus on analyzing phonesthemic information contained in word embedding spaces (abramova2013automatic; abramova-fernandez-2016-questioning).
-Recent studies demonstrate that the effect of non-arbitrary integration of linguistic form and meaning is substantiated in text-only LLMs (miyakawa-etal-2024-llms; marklova2025iconicitylargelanguagemodels), vision and image generation models (loakman-etal-2024-ears; alper2024kikiboubasoundsymbolism; shinto2024analyzing; iida2024investigating), and audio-visual models (tseng2024measuring), yet they are not extensive enough in scope to include exhaustive and multilingual mimetic word data or diverse semantic feature analysis.
+Iconicity tasks for language models have been used as a means of assessing whether the models have human-like phonetic intuition (Cai et al.
+2024; Duan et al.
+2024).
+Early studies focus on analyzing phonesthemic information contained in word embedding spaces (Abramova et al.
+2013; Abramova and Fernández 2016).
+Recent studies demonstrate that the effect of non-arbitrary integration of linguistic form and meaning is substantiated in text-only LLMs (Miyakawa et al.
+2024; Marklová et al.
+2025), vision and image generation models (Loakman et al.
+2024; Alper and Averbuch-Elor 2024; Shinto and Iizuka 2024; Iida and Funakura 2024), and audio-visual models (Tseng et al.
+2024), yet they are not extensive enough in scope to include exhaustive and multilingual mimetic word data or diverse semantic feature analysis.
 2.3 Multimodal Interpretability
-Recently, the field of model interpretability has witnessed significant growth (elhage2021mathematical; zou2025representationengineeringtopdownapproach; wang2023interpretability).
-Motivated by the rapid progress of MLLMs (openai2024gpt4ocard; xu2025qwen25omnitechnicalreport), there has been increasing interest in multimodal interpretability (lin2025surveymechanisticinterpretabilitymultimodal).
-For instance, neo2025towards ablates a subset of visual tokens to observe the resulting differences in model output, while nikankin2025taskdifferentcircuitsdisentangling demonstrates that models employ distinct processing circuits depending on the input modality.
+Recently, the field of model interpretability has witnessed significant growth (Elhage et al.
+2021; Zou et al.
+2025; Wang et al.
+2023).
+Motivated by the rapid progress of MLLMs (OpenAI et al.
+2024; Xu et al.
+2025), there has been increasing interest in multimodal interpretability (Lin et al.
+2025).
+For instance, Neo et al.
+(2025) ablates a subset of visual tokens to observe the resulting differences in model output, while Nikankin et al.
+(2025) demonstrates that models employ distinct processing circuits depending on the input modality.
 Despite these advances, prior work has predominantly focused on the visual modality, with less exploration of audio-based interpretability.
-yang2025audiolenscloserlookauditory investigates how models handle auditory input, but their analysis is limited to simple settings where the audio simply serves as a direct vocalization of textual data.
+Yang et al.
+(2025b) investigates how models handle auditory input, but their analysis is limited to simple settings where the audio simply serves as a direct vocalization of textual data.
 In this work, we examine audio interpretability through the lens of sound symbolism, highlighting how phoneme-level features encode meaning in ways that are intrinsically tied to the auditory modality.
@@
 Table 1: Data distribution of LEX-ICON across natural (8,052 words from English, French, Japanese, and Korean) and constructed (2,930 words) groups.
-Japanese and Korean are known for their rich mimetic words (hamano1986sound; kwon2018iconicity), which contributes to their high proportion.
+Japanese and Korean are known for their rich mimetic words (Hamano 1986; Kwon 2018), which contributes to their high proportion.
 Dimension Natural Constructed Total
@@
 Table 2: Semantic dimension distribution of pseudo ground truth data by word group.
-We adopt 25 semantic dimension criteria by sidhu2022higher to annotate LEX-ICON.
+We adopt 25 semantic dimension criteria by Sidhu et al.
+(2022) to annotate LEX-ICON.
 Six dimensions from the constructed word group are excluded by removing close-to-neutral data points.
@@
 3.1 Semantic Dimension
-We employ the semantic dimension methodology, a concept originated from ``semantic differential'' by osgood1957measurement, which consists of two semantic feature adjectives located at opposite extremes, such as ``big'' and ``small.'' By applying this methodology to sound symbolism experiments, we can simultaneously and precisely measure the multifactorial meanings contained in a single word.
-We annotate up to 25 pairs of predefined semantic features as per sidhu2022higher for each word in the LEX-ICON, which is the most diverse scale in terms of studies on LLMs' iconicity.
+We employ the semantic dimension methodology, a concept originated from ``semantic differential'' by Osgood et al.
+(1957), which consists of two semantic feature adjectives located at opposite extremes, such as ``big'' and ``small.'' By applying this methodology to sound symbolism experiments, we can simultaneously and precisely measure the multifactorial meanings contained in a single word.
+We annotate up to 25 pairs of predefined semantic features as per Sidhu et al.
+(2022) for each word in the LEX-ICON, which is the most diverse scale in terms of studies on LLMs' iconicity.
 3.2 Natural Mimetic Words
@@
 We manually collect 8,052 mimetic words and definition data consisting of onomatopoeias and ideophones in English, French, Japanese, and Korean from specialized mimetic word lexicons and authoritative dictionaries for each language.
-We then extract the most representative definitions of these mimetic words from dictionaries such as the Oxford English Dictionary (dictionary1989oxford), Le Petit Robert (lepetitrobert2023), Nihon Kokugo Daijiten (9784095210216), and the Standard Korean Language Dictionary (nkl2025).
+We then extract the most representative definitions of these mimetic words from dictionaries such as the Oxford English Dictionary (Simpson 1989), Le Petit Robert (Alain Rey 2022), Nihon Kokugo Daijiten (Shogakukan 2006), and the Standard Korean Language Dictionary (National Institute of Korean Language 2025).
 For further source information, see the Appendix.
@@
 To observe the effect of trained token memorization and modality change, we create three types of word form that contain textual and auditory input: (1) original text, (2) IPA-converted text with phoneme-level spacing, (3) audio waveform converted using text-to-speech (TTS) software.
-We perform IPA conversion with the Epitran package (Mortensen-et-al:2018), and obtain the TTS dataset using Google Text-to-Speech (google_tts_api_2025) for English, French, and Japanese; and MeloTTS (zhao2024melo) for Korean.
+We perform IPA conversion with the Epitran package (Mortensen et al.
+7–12), and obtain the TTS dataset using Google Text-to-Speech (Google 2025) for English, French, and Japanese; and MeloTTS (Zhao et al.
+2023) for Korean.
 Large-Scale Annotation Process.
-To effectively generate a large-scale ground truth data, all natural language words and their definitions in the LEX-ICON are given four LLMs: GPT-4.1 (openai2025_gpt41), Qwen3-32B (yang2025qwen3technicalreport) [^2] Non-reasoning mode., Gemma-3-27B (gemmateam2025gemma3technicalreport), and Gemini-2.5-flash (comanici2025gemini25pushingfrontier).
+To effectively generate a large-scale ground truth data, all natural language words and their definitions in the LEX-ICON are given four LLMs: GPT-4.1 (OpenAI 2025), Qwen3-32B (Yang et al.
+2025a) [^2] Non-reasoning mode., Gemma-3-27B (Team et al.
+2025b), and Gemini-2.5-flash (Comanici et al.
+2025).
 The models are prompted to annotate each semantic dimension of a word with one of a feature pair or neutral labels (e.g., selecting one of the ``exciting'', ``calming'', or ``neither'' option) for each word.
@@
 We use four vowels from two categories: front vowels (/i/, /ej/) and back vowels (/A/, /ow/).
-After filtering against existing entries in the IPA-dict database (ipa_dict) across four languages (US and UK English, French, Japanese, and Korean), we obtain 3,108 novel pseudo-words.
+After filtering against existing entries in the IPA-dict database (Doherty 2016) across four languages (US and UK English, French, Japanese, and Korean), we obtain 3,108 novel pseudo-words.
 We convert IPA symbols to English alphabet combinations for TTS compatibility and remove incorrectly pronounced words by Google TTS, yielding 2,930 final words.
 Semantic Dimension Annotation.
-We utilize the empirical coefficients from sidhu2022higher, who quantified associations between phoneme categories and semantic dimensions through human rating experiments.
+We utilize the empirical coefficients from Sidhu et al.
+(2022), who quantified associations between phoneme categories and semantic dimensions through human rating experiments.
 These coefficients represent how much phoneme categories deviate from the overall mean rating across 25 semantic dimensions.
@@
 4.1 Experimental Settings
-We employ MLLMs that officially support simultaneous text and audio inputs: Qwen2.5-Omni (xu2025qwen25omnitechnicalreport) [^3] Qwen2.5-Omni-3B and Qwen2.5-Omni-7B., Gemini-2.5-flash (geminiteam2025geminifamilyhighlycapable), and GPT-4o (openai2024gpt4ocard)444gpt-4o for text-only input, gpt-4o-audio-preview for audio-enabled input..
+We employ MLLMs that officially support simultaneous text and audio inputs: Qwen2.5-Omni (Xu et al.
+2025) [^3] Qwen2.5-Omni-3B and Qwen2.5-Omni-7B., Gemini-2.5-flash (Team et al.
+2025a), and GPT-4o (OpenAI et al.
+2024)444gpt-4o for text-only input, gpt-4o-audio-preview for audio-enabled input..
 All experimental models cover the four languages and orthographies present in LEX-ICON.
@@
 Notably, the models' strong performance on the sharp vs.
-round dimension aligns with well-known cognitive linguistic experiments like the ``bouba-kiki'' effect (ramachandran2001synaesthesia).
+round dimension aligns with well-known cognitive linguistic experiments like the ``bouba-kiki'' effect (Ramachandran and Hubbard 2001).
 Human-like Iconicity.
@@
 For dimensions where acoustic features are theoretically central, such as size distinctions (big vs.
-small) that correlate with formant frequencies (knoeferle2017drives) and speed distinctions (fast vs.
-slow) that relate to consonant voicing duration (saji2013cross), the MLLMs show a pattern of enhanced performance when processing constructed words in audio format, consistent with the theoretical predictions.
+small) that correlate with formant frequencies (Knoeferle et al.
+2017) and speed distinctions (fast vs.
+slow) that relate to consonant voicing duration (Saji et al.
+2013), the MLLMs show a pattern of enhanced performance when processing constructed words in audio format, consistent with the theoretical predictions.
 Conversely, for dimensions where non-acoustic mechanisms are proposed, such as shape associations (sharp vs.
-round) based on lip rounding gestures (imai2025does) and valence associations (beautiful vs.
+round) based on lip rounding gestures (Imai et al.
+2025) and valence associations (beautiful vs.
 ugly, happy vs.
-sad) affected by articulatory properties (korner2022articulation), the models exhibit stronger reliance on textural representations.
+sad) affected by articulatory properties (Körner and Rummer 2022), the models exhibit stronger reliance on textural representations.
 Figure 5: Advantage scores (macro-F1 differences) of audio inputs over the original text inputs by word group.
@@
 For each IPA and semantic dimension, the model tends to attend more to semantic features that exhibit stronger associations with the given phonemes in linguistics.
-For instance, the model mostly associates sharp semantic feature with /p/ and /k/, round with /m/ and /n/, big with /A/, and small with /i/ (kohler1967gestalt; parise2012audiovisual).
+For instance, the model mostly associates sharp semantic feature with /p/ and /k/, round with /m/ and /n/, big with /A/, and small with /i/ (Köhler 1967; Parise and Spence 2012).
 5.1 Experimental settings
@@
 When an IPA symbol spans multiple text tokens, the procedure first sums the attention scores across those tokens before fraction normalization.
-For audio inputs, we employ Montreal Forced Aligner (mcauliffe2025montreal) to segment each audio waveform into phonemes over time.
+For audio inputs, we employ Montreal Forced Aligner (McAuliffe and others 2025) to segment each audio waveform into phonemes over time.
 The phoneme sequence is then aligned with the model’s 40 ms sampling period, with consecutive occurrences of the same phoneme treated as a single IPA symbol.
@@
 The patterns for constructed words in the IPA modality closely mirror prior findings in sound symbolism research.
-For example, phonemes such as /p/ and /k/ exhibit elevated attention under the sharp feature, whereas /m/ and /n/ associate with the round feature (kohler1967gestalt).
+For example, phonemes such as /p/ and /k/ exhibit elevated attention under the sharp feature, whereas /m/ and /n/ associate with the round feature (Köhler 1967).
 Figure 7: Attention fraction scores for the constructed word group, averaged across semantic dimensions.
@@
 We manually extract headwords and definitions from published onomatopoeia dictionaries for each language to determine whether a word belongs to the mimetic word category.
-For Korean, the word definition guideline of the NIKL_StandardKoreanLanguageDictionaryCompilationGuideline2_2000 is used to extract mimetic words and definitions.
+For Korean, the word definition guideline of the National Institute of Korean Language (2000) is used to extract mimetic words and definitions.
 Semantic definitions that need to be supplemented are obtained from authoritative dictionaries for each language.
@@
 Language References
-English Rhyme over Reason: Phonological Motivation in English (benczes2019rhyme)
-Oxford English Dictionary, 2nd Edition (dictionary1989oxford)
-French Dictionnaire des Onomatopées (enckell2003dictionnaire)
-Le Petit Robert de la Langue Française (lepetitrobert2023)
-Trésor de la Langue Française (letresor1994)
-Japanese Kurashi no Kotoba Gion Gitaigo Jiten (1130282271732203008)
-Seisenban Nihon Kokugo Daijiten (9784095210216)
-Kojien, 7th Edition (kojien7)
-Digital Daijisen (digitaldaijisen)
-Daijirin, 4th Edition (matsumura2019daijirin)
-Korean Standard Korean Language Dictionary (nkl2025)
+English Rhyme over Reason: Phonological Motivation in English (Benczes 2019)
+Oxford English Dictionary, 2nd Edition (Simpson 1989)
+French Dictionnaire des Onomatopées (Enckell and Rézeau 2003)
+Le Petit Robert de la Langue Française (Alain Rey 2022)
+Trésor de la Langue Française (Centre national de la recherche scientifique and Institut national de la langue française 1994)
+Japanese Kurashi no Kotoba Gion Gitaigo Jiten (Yamaguchi 2003)
+Seisenban Nihon Kokugo Daijiten (Shogakukan 2006)
+Kojien, 7th Edition (Shinmura 2018)
+Digital Daijisen (Shogakukan 2025)
+Daijirin, 4th Edition (Matsumura 2019)
+Korean Standard Korean Language Dictionary (National Institute of Korean Language 2025)
 Table 4: Reference materials for natural mimetic word collection by language.
@@
 A.2 Semantic Dimension
-Table 6 shows the extreme feature pairs of 25 semantic dimensions suggested by sidhu2022higher, which have been utilized for our semantic dimension prediction experiments.
+Table 6 shows the extreme feature pairs of 25 semantic dimensions suggested by Sidhu et al. (2022), which have been utilized for our semantic dimension prediction experiments.
 After the automatic annotation process for natural words is complete, we utilize labels that are agreed upon by all model families and not annotated as ``neither'' as pseudo ground truth.
@@
 dangerous safe
-Table 6: The 25 semantic dimension list by sidhu2022higher.
+Table 6: The 25 semantic dimension list by Sidhu et al. (2022).
 A.3 An Example Entry from LEX-ICON
@@
   • Hardware: four NVIDIA RTX 4090 GPUs for 27B and 32B models, one NVIDIA RTX 4090 GPU for 3B and 7B models, and API calls for other proprietary models.
-  • Software: PyTorch (paszke2017automatic) and vLLM (kwon2023efficient) libraries.
+  • Software: PyTorch (Paszke et al. 2017) and vLLM (Kwon et al. 2023) libraries.
   • Random seed: 42 (for reproducibility, only if needed).
@@
 In this work, we utilize the following language models for data annotation, experimentation, and analysis.
-  • gpt-4o (openai2024gpt4ocard)
-  • gpt-4o-audio-preview (openai2024gpt4ocard)
-  • gpt-4.1 (openai2025_gpt41)
-  • gemini-2.5-flash (comanici2025gemini25pushingfrontier)
-  • gemma-3-27b-it (gemmateam2025gemma3technicalreport)
-  • Qwen3-32B (yang2025qwen3technicalreport)
-  • Qwen2.5-Omni-3B (xu2025qwen25omnitechnicalreport)
-  • Qwen2.5-Omni-7B (xu2025qwen25omnitechnicalreport)
+  • gpt-4o (OpenAI et al. 2024)
+  • gpt-4o-audio-preview (OpenAI et al. 2024)
+  • gpt-4.1 (OpenAI 2025)
+  • gemini-2.5-flash (Comanici et al. 2025)
+  • gemma-3-27b-it (Team et al. 2025b)
+  • Qwen3-32B (Yang et al. 2025a)
+  • Qwen2.5-Omni-3B (Xu et al. 2025)
+  • Qwen2.5-Omni-7B (Xu et al. 2025)
 Appendix C Token-IPA Symbol Alignment
@@
 C.1 Montreal Forced Alignment (MFA) Pipeline
-We employ Montreal Forced Alignment (mcauliffe2025montreal) to obtain precise phoneme-level segmentation of audio waveforms. We utilize pre-trained acoustic models for each language:
-  • English (US) ARPA acoustic model v3.0.0 (mfa_english_us_arpa_acoustic_2024)
-  • French MFA acoustic model v3.0.0 (mfa_french_mfa_acoustic_2024)
-  • Japanese MFA acoustic model v3.0.0 (mfa_japanese_mfa_acoustic_2024)
-  • Korean MFA acoustic model v3.0.0 (mfa_korean_mfa_acoustic_2024)
+We employ Montreal Forced Alignment (McAuliffe and others 2025) to obtain precise phoneme-level segmentation of audio waveforms. We utilize pre-trained acoustic models for each language:
+  • English (US) ARPA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024a)
+  • French MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024b)
+  • Japanese MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024c)
+  • Korean MFA acoustic model v3.0.0 (McAuliffe and Sonderegger 2024d)
 The alignment process generates Textgrid files containing phoneme boundaries with millisecond precision.
@@
 To ensure the reliability of the semantic dimension selection and the automatic annotation of the pseudo ground truth data in LEX-ICON, we conduct a human evaluation using 152 questions randomly sampled for each participant from the entire dataset.
-10 graduate students without expertise in linguistics participate in this experiment, which is performed via the Label Studio (LabelStudio).
+10 graduate students without expertise in linguistics participate in this experiment, which is performed via the Label Studio (Tkachenko et al.
+2020).
 The experiment utilize questions with the same prompt structure as the semantic dimension prediction test that was provided to LLMs.